### v1

Abstract
Deduplication is a major focus for assembling and curating training datasets for large language models (LLM) – detecting and eliminating additional instances of the same content – in large collections of technical documents. Unrestrained, duplicates in the training dataset increase training costs and lead to undesirable properties such as memorization in trained models or cheating on evaluation. Contemporary approaches to document-level deduplication are often extremely expensive in both runtime and memory. We propose LSHBloom, an extension to MinhashLSH, which replaces the expensive LSHIndex with lightweight Bloom filters. LSHBloom demonstrates the same deduplication performance as MinhashLSH with only a marginal increase in false positives (as low as 1e-5 in our experiments); demonstrates competitive runtime (270% faster than MinhashLSH on peS2o); and, crucially, uses just 0.6% of the disk space required by MinhashLSH to deduplicate peS2o. We demonstrate that this space advantage scales with increased dataset size—at the extreme scale of several billion documents, LSHBloom promises a 250% speedup and a 54$\times$ space advantage over traditional MinHashLSH scaling deduplication of text datasets to many billions of documents.
marginparsep has been altered.
topmargin has been altered.
marginparwidth has been altered.
marginparpush has been altered.
The page layout violates the ICML style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
LSHBloom: Memory-efficient, Extreme-scale Document Deduplication
Anonymous Authors1
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Scholarly publications and related literature represent a tremendous body of often carefully encoded human knowledge. Google Scholar is estimated to index some 400 million documents Gusenbauer (2019), suggesting that the total number of unique documents worldwide numbers at least in the hundreds of millions. And a growing number of digital repositories mean that researchers can today access millions of such documents with ease, and with effort, tens of millions or more.
Developments in large language models (LLMs) suggest that such corpora can be used to develop powerful agents able to assist with technical work in a variety of fields Zhang et al. (2024). However, seizing this opportunity at the largest scale requires addressing challenging deduplication problems. Any large corpus of technical documents assembled from diverse locations will include multiple copies of many documents: not only identical copies but also copies that vary, subtly or otherwise, for reasons such as different publication formats, the use of different parsers, copyediting, and versioning. As is well known, duplication in training datasets can lead to a host of undesirable behaviors, from excessive training times to predictive biases and training data memorization Lee et al. (2021) in trained models. Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets Chandran et al. (2024), illegitimately increasing perceived model performance by 8–11% Dekoninck et al. (2024).
These considerations motivate the problem that we tackle in this paper, namely the accurate and efficient deduplication of large numbers (10s to 100s of millions) of scholarly and technical documents. We approach this problem from two perspectives. First, we demonstrate via careful evaluation that existing state-of-the-art approaches to the deduplication of technical documents suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs. Due to the frequent loss of metadata and variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method Broder et al. (1998). The infeasibility of pairwise comparisons in large datasets motivated the development of the MinHashLSH method Leskovec et al. (2014), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents. However, the associated MinHashLSH index can grow extremely large. For example, applying MinHashLSH to the peS2o dataset Soldaini & Lo (2023) of just 8M full-text and 31M abstracts takes more than 14 hours on a single 32-core node of the Polaris supercomputer and requires more than 300GB of disk space (see Section 5.3.1).
Second, we introduce a new method, LSHBloom, that improves significantly on the state of the art in terms of the computational costs required to achieve a desired level of deduplication performance. The key insight here is that the process that MinHashLSH employs to find matches between MinHash signatures can be approximated, with a modest and controllable loss of accuracy, by using Bloom filters. This use of Bloom filters enables LSHBloom to run almost three times faster than MinHashLSH by interfacing with inexpensive bit arrays rather than inserting and querying against the comparatively slow tree data structures that comprise the traditional LSHIndex. Most importantly, this use of Bloom filters means that LSHBloom requires just 0.6% of the disk space consumed by the traditional MinHashLSH index to deduplicate peS2o Soldaini & Lo (2023), a dataset of 39 million documents. We demonstrate that this space advantage remains as the number of documents increases and makes it feasible for LSHBloom to process several billion documents with space savings of 54$\times$ relative to MinHashLSH without sacrificing deduplication quality.
2 Background
We first provide background information on the deduplication problem and describe the current state-of-the-art for large-scale text deduplication pipelines.
2.1 Defining the Deduplication Problem
Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus. These methods can be applied at different resolutions (e.g., to documents, paragraphs, sentences, or even tokens) and scopes (e.g., across a corpus or within a single document). Deduplication methods may also differ in how they define duplication (e.g., exact matches, n-gram overlap, “fuzzy”). In evaluating a deduplication method for a particular application, we are concerned with its precision, recall, and computational performance. Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces the size of the training dataset and limits an LLM’s ability to learn the underlying distribution of the text data. Low recall, i.e., excessive false negatives, admits a great degree of duplicate content in our dataset, with negative effects on learning as well as potentially inducing degenerate behavior such as memorization. High computational performance is critical due to the extreme size of modern LLM datasets. For example, peS2o Soldaini & Lo (2023) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature Gusenbauer (2019), DOLMA Soldaini et al. (2024) contains roughly 5B documents, and RedPajama Computer (2023) has just over 100B documents—scales at which naive pairwise comparisons become infeasible. Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
Here, we focus on methods for document-level deduplication. Suppose an operator $S(d_{i},d_{j}):D\times D\rightarrow[0,1]$ that takes as input two documents and outputs a similarity score between 0 and 1. We define our task as follows: Given a corpus of $N$ documents, $D_{N}=$$d_{1}$, …, $d_{N}$}, and a similarity threshold, $T\in[0,1]$, identify every document $d$ such that $\exists d^{\prime}\in D_{N}$ such that $d^{\prime}\neq d$ and $S(d,d^{\prime})\geq T$.
Just how to define duplication is a subtle question. For example, one could stipulate simply that two documents must be byte-for-byte identical: equivalent to setting $S$ to measure byte overlap and $T$ = 1.0 above. However, this definition may exclude many document pairs that we would like to consider duplicates due to great similarity in their token distributions or semantic content, as are encountered frequently in practical text ingestion pipelines. For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions. When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and various format-specific text parsing pipelines tend to induce small differences even when applied to the same underlying document. Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language. PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
For these reasons, many deduplication techniques choose to take a “fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs. One way to do this is to alter one’s choice of similarity metric to encapsulate a broader semantic notion of similarity. This is common in so-called semantic deduplication techniques such as SemDeDup Abbas et al. (2023), which employ embeddings from pre-trained language models to identify semantically similar text data. Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents. It is more efficient to instead establish a similarity metric based on the raw text content of each document.
2.2 Jaccard Similarity and MinHash Approximation
Popular techniques such as those used to build the DOLMA Soldaini et al. (2024) and DataComp-LM Li et al. (2024) datasets measure n-gram overlap between documents to identify duplicates. Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
A text document can be viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets. For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming. The MinHashing procedure Broder et al. (1998) provides a means of efficiently approximating the Jaccard similarity of two documents. Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures. The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets Broder et al. (1998); Leskovec et al. (2014). Therefore, we can repeatedly apply random permutations to the two sets’ elements and compare them for equality to obtain an estimate of the Jaccard similarity. Our estimate grows more accurate as we increase our sample size.
2.3 MinHashLSH for Scalable MinHash
The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed computed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document. However, the need to compare every pair of documents for similarity during deduplication still introduces a quadratic cost when comparing all documents in a corpus.
Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs. MinHashLSH Leskovec et al. (2014) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents. Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document. Then, given a desired Jaccard similarity threshold, $T$, one can determine an optimal band size, $b$, with which to group the rows of the MinHash signature matrix. This results in $b$ groupings of rows (bands) of size $r$. With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$. This enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity. This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication. However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 23 TB for 5B documents. Additionally, inserting and querying these tree-structured indices constitutes a significant portion (>80%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1). Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
Figure 1: Breakdown of wall clock time on 5% of peS2o for a conventional LSH algorithm and our LSHBloom method
2.4 Bloom Filter
The Bloom filter Bloom (1970) is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error. That is, the Bloom filter may admit false positives but not false negatives.
A Bloom filter consists of a bit array of size $m$, with all bits initially set to 0. $k$ independent hash functions are defined, each of which hashes an element to an integer in the range 1 to $m$. To insert an element into the Bloom filter, it is passed through the $k$ hash functions, and the resulting $k$ positions in the bit array are set to 1. We can then check if an element is in the set by hashing using the same $k$ hash functions. If all $k$ of the corresponding positions in the bit array are 1, the filter returns a positive. If any of those $k$ bits is 0, the element is not contained in the set.
Bloom filters have been used extensively in distributed computing for a variety of purposes Broder & Mitzenmacher (2004); Chervenak et al. (2008); Tarkoma et al. (2011).
3 Related Work
Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs. LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data. Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 Touvron et al. (2023b), Claude, Gemini Gemini Team (2024a; b), and GPT versions 3 and later OpenAI (2024). Others, such as Palm2 Anil et al. (2023) and Gemma Gemma Team (2024); Gemini Team (2024b), provide only limited details. Others, including several leading models, provide extensive details about how data for their model was dededuplicated including Olmo’s Dolma dataset Soldaini et al. (2024) as well as its successor DataComp-LM Li et al. (2024), Llama 1 Touvron et al. (2023a), Llama 3 Dubey et al. (2024), and OPT Zhang et al. (2022). For example, Llama 3 Dubey et al. (2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data. Phi uses an approach described by Gunasekar et al. (2023) and elaborated on in their Phi 2.0 report Javaheripi (2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation.
The most common methods used across these model sources include metadata-based approaches, MinHashLSH, CCNet, and n-gram-based methods. Metadata-based approaches (e.g., exact URL matching, DOIs) are by far the simplest approach; however, they are applicable only in cases in which a reliable source of metadata exists, such as those provided by a single publisher. Unfortunately, for our target datasets this metadata is often unavailable or is otherwise untrustworthy. This can occur, for example, when: 1) a document appears in two or more collections with different metadata schemes (e.g., ArXiV vs. ACM, different GitHub repositories); 2) sections/paragraphs of text are copied, as in extended journal versions of computer science papers; or 3) metadata is not curated (e.g., an ArXiV entry may not be updated to reflect a subsequent publication).
CCNet Wenzek et al. (2020) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content. CCNet performs dataset sharding, deduplication, language identification, and quality filtering. CCNet’s deduplication technique begins by preprocessing text so that all characters are lowercase and eliminates special Unicode characters. CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates. CCNet can optionally compare to a single document, a subset of documents, or all documents.
Dolma Soldaini et al. (2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma). An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter. If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate. By default, $T$ = 1.0 in their configuration.
DataComp-LM Li et al. (2024) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter. They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer Ye et al. (2023). For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams. If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document. If the proportion of duplicated n-grams in the entire document exceeds the user-defined threshold, then the document is removed from the dataset. They employ the same threshold parameters for paragraph-level and document-level deduplication.
4 The LSHBloom Method
4.1 Overview
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for hash collisions. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of Bender et al. (2022).
4.2 Insertion
Given a Jaccard similarity threshold $T\in$ [0, 1], MinHashLSH traditionally groups signatures into a series of $b$ bands of size $r$ such that an exact match in one band for two columns of the MinHash signature matrix denotes a duplicate document. That is, with high probability, the two documents with a matching band have Jaccard similarity greater than or equal to the threshold. To avoid storing the signatures in full, as is required by MinHashLSH, we assign each band of the minhash signature matrix its own Bloom filter. Thus we have in total $b$ Bloom filters, and a collision in any one indicates a duplicate document.
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers onto a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function, with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values Carter & Wegman (1977). Then, we can hash a vector $\bar{x}$ of $r$ integers as follows:
$h(\bar{x})=\left(\sum_{i=1}^{r}h_{i}(x_{i})\right)\mathrm{mod}\ {N}$
This approach enables us to map each band to a single integer value with known collision probability. To insert a document into the index, we simply insert each of the $b$ final integer values into each of the $b$ Bloom filters corresponding to the $b$ bands in the signature matrix.
4.3 Querying
Querying a document against the index is a similar process to insertion. We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in Section 4.2. If there is a collision for any band of signatures, we mark that document as a duplicate. This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters. However, the false positive probability is configurable and can be set to be arbitrarily low (in our experiments, we set the effective false positive rate as low as 1e-5).
4.4 Advantages
We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams. Since this representation stems from the k-shingling of documents and many practical scenarios may admit an extremely large universe of such k-shingles (e.g., if we allow our set of documents to include a variety of encodings, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the range of MinHash hash values or increasing the accuracy of MinHash estimates of the Jaccard similarity. This can be done by increasing the number of permutations applied to prevent as much information loss as possible via compression of the text representation. A scalable deduplication algorithm should be amenable to such considerations. LSHBloom allows for the space-efficient representation of extremely large deduplication indices. This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with the same number of bits. If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits Bender et al. (2022).
[…]
As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing. For these parameter settings, MinHashLSH creates nine bands for the MinHash signature matrix, meaning that we require nine Bloom filters for our index. For these settings, setting a conservative $p$effective = 1e-10 for a dataset of 10B documents would only require 590 GB of space. Meanwhile, traditional MinHashLSH may require 46 TB to store the equivalent LSHIndex—exemplifying the scalability of our technique.
While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration. The additional false positive probability scales with the number of documents we intend to insert and scales inversely with the size of each filter. Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter. However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.
5 Evaluation
Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance. As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets. For correctness, we compare LSHBloom against state-of-the-art deduplication methods when applied to a dataset of 25K PDFs containing labeled duplicates. Without correctness, the system may either under-identify duplicates, resulting in decreased downstream inference quality; or over-identify duplicates, unnecessarily decreasing the volume of the training dataset. For performance, we consider the storage utilization and runtime of each method on a larger collection of PDFs, peS2o Soldaini & Lo (2023). Without performance, it may not be possible to process large datasets on available hardware resources.
We conduct a comprehensive analysis of various deduplication methods. Specifically, as each method includes a set of parameters whose values affect both their correctness and resource usage, we evaluate each using a range of parameter settings. Understanding these relationships is crucial to making the best use of available resources.
In the following subsections, we describe the details of our evaluation beginning with the evaluation methodology, then proceeding to correctness, performance, and finally the tradeoffs between performance and correctness.
5.1 Experimental Methodology
In this subsection, we describe our experimental methodology including testbed hardware, state-of-the-art methods for comparison, our evaluation metrics, and datasets.
5.1.1 Testbed Hardware
We ran our experiments on the Polaris supercomputer, a 560-node HPE Apollo 6500 Gen 10+-based system. Each node is equipped with a 32-core AMD Zen 3 2.8 GHz processor, with 512 GB of DDR4 RAM, and 3.2 TB of local SSD: representative of nodes found in large-scale high-performance computing facilities used for deduplication tasks.
5.1.2 Baselines
We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM. We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language. This also enabled us to use the same Bloom filter implementation for each technique. Below, we describe any changes or configurations to the methods.
MinhashLSH Broder (1997) (lsh): We configure the traditional MinhashLSH algorithm as described in Section 2.3. DOLMA: We apply both standard dolma and dolma-ngram methods to document-level deduplication. Specifically, for dolma we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$. For dolma-ngram, we split each document into a set of n-grams with a given size, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams that are duplicated within that document to the overlap threshold, $T$. Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication. CCNet originally performs deduplication at the paragraph level; however, here, we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1]. For a fair comparison, we compare documents to all other documents—just as MinhashLSH does. When evaluating memory usage, we apply all CCNet optimizations, such as the use of memory-efficient hash sets. Additionally, we augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a more appropriate comparison during our scaling tests. DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
When considering the correctness of techniques that operate on n-grams and use Bloom filters, we must give an expected n-gram count in the corpus; choosing this count accurately is crucial to make a fair comparison. Because counting the number of n-grams of a given size in a corpus is far more computationally expensive than counting the number of documents (as is required for our technique), we compute an estimate of the n-gram count in a corpus. To do so we randomly sample $N$ = 1000 documents from the corpus, compute the mean n-gram count from our sample, and multiply the mean n-gram count by the number of total documents in that corpus to obtain an estimate of the number of n-grams in the entire corpus. We parameterize Bloom filters for n-gram techniques using this estimate.
5.1.3 Metrics
Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents. We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance. Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives. Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization. To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using the F1 score:
$F1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$
Performance: To assess the resource costs of each technique we measure wall clock time and disk space usage. We measure wall clock time using /user/bin/time and assess each method’s disk space requirements using Python’s standard os.path.getsize method.
5.1.4 Constructing the correctness evaluation dataset
To permit the rigorous evaluation of de-duplication methods, we created an evaluation dataset By acquiring 24,000 scientific articles (4156 scraped from HTML and 20,844 parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata. We then duplicated 1000 of those documents for a total of 25,000. We then translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF Smith (2024), Nougat Blecher et al. (2023), Marker Paruchuri (2024), Grobid Lopez (2009), or Tesseract Smith (2007), each used with roughly the same frequency—for a total of 24,956 HTML or parsed PDF documents (as 44 failed to parse). Finally, we transformed 419 of these documents by removing a randomly selected subset of its tokens
5.1.5 Choice of values for Tuning Parameters
[…]
It is possible that one could properly tune these deduplication methods further using incremental changes in their parameter settings. However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance. Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs. Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations. MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget. This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see Section 5.3.1).
(a) MinHashLSH
(b) LSHBloom
Figure 2: F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
5.2 Deduplication Correctness
5.2.1 LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 2(a) and 2(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate. However, there is a clear saturation point at 128 permutation functions, after which we observe only marginal performance increases. This is particularly relevant for LSHBloom, for which increasing the number of permutation functions blindly can lead to slight degradations in performance—due to the introduction of more bands in the signature matrix by the LSH function, requiring more Bloom filters, and therefore leading to an increase in the false positive rate. This can be avoided, however, by jointly tuning the false positive rate and the number of permutation functions. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.6 yields the best performance, as more stringent thresholds lead to a diminishing F1 score.
(a) DOLMA
(b) DataComp-LM
Figure 3: F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
5.2.2 Paragraph Level and N-Gram Methods
Figure 4 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. N-gram methods such as DOLMA-ngram and CCNet do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Figures 3(a) and 3 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, both paragraph and n-gram methods demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates. In contrast, n-gram methods yield many false negatives at high thresholds and yield many false positives at low thresholds.
Figure 4: F1 score vs. threshold for paragraph-level techniques.
5.3 Deduplication Resource Usage
(a) Quality vs. Wall Clock Time
(b) Quality vs. Disk
Figure 5: Pareto plots for F1 score vs. resource usage.
As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents. Here we work with the peS2o dataset Soldaini & Lo (2023) of roughly 8M full-text and 31M abstracts. We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter). We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost. DataComp-LM, for example, takes one hour just to deduplicate 25,000 examples in our previous benchmark in Section 5.1.4. Even processing just 10% of peS2o using this technique would require 160 hours by extrapolation. We discuss the results for other methods below.
Figure 6: Disk size of index for different methods, as measured for peS2o subsets of increasing size.
Figure 7: Wall clock times for different methods, as measured for peS2o subsets of increasing size.
5.3.1 Scaling of Resource Usage
In this section, we study how resource usage grows with the number of documents. In these experiments, we set any threshold parameters to 0.8, the number of random permutations for our LSH techniques to 128, and $p$ = 1e-5 for all Bloom filters. Ideally, we would have informed these parameter settings based on the optimal settings from Section 5.1.4; however, scaling experiments on peS2o regularly took over a day to complete, limiting our ability to tune parameters and necessitating heuristic choices for parameter settings as we evaluated deduplication quality in parallel.
Figure 7 shows the wall clock time to deduplicate various subsets of peS2o with each method. Notably, MinHashLSH incurs the largest runtime while paragraph-level techniques such as DOLMA and CCNet are relatively fast. However, DOLMA and CCNet achieve this speed with significantly lower deduplication performance: see Section 5.1.4. Meanwhile, LSHBloom, which has comparable deduplication performance to MinHashLSH, is much faster. To deduplicate all of peS2o, LSHBloom takes just five hours, whereas MinHashLSH takes 14 hours—almost a 3$\times$ speedup. Figure 1 shows the breakdown of this time for MinHashLSH and LSHBloom; we can see that, as expected, the time to compute MinHashes is the same; however, the insertion and query latencies are significantly reduced in the Bloom filter-based index compared to the traditional LSHIndex.
Figure 6 shows the disk usage for each method. We again see a similar trend. The paragraph-level techniques maintain only a single Bloom filter which leads to extremely low disk usage. Meanwhile, MinHashLSH’s index grows steadily in size as the number of documents increases. LSHBloom’s index also grows linearly in the number of documents but at a reduced rate. When processing all of peS2o, LSHBloom requires just 1 GB vs. the 180 GB taken by MinHashLSH—a 180$\times$ reduction.
5.3.2 Estimated Resource Usage at Scale
Figure 8: Predicted wall clock times vs. number of documents (in Billions) for different methods
[…]
Table 1: Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
Technique Bloom filter false Index disk
positive overhead usage (GB)
MinHashLSH 23,008.04
LSHBloom 1e-5 160.51
LSHBloom 1e-10 295.30
LSHBloom 1e-15 425.35
Table 2: Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
Technique Bloom filter false Index disk
positive overhead usage (TB)
MinHashLSH 460.16
LSHBloom 1e-5 3.21
LSHBloom 1e-10 5.91
LSHBloom 1e-15 8.51
5.4 Performance vs. Quality Tradeoffs at Scale
Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance. Figures 5(a) and 5(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. Additionally, LSHBloom displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown in Section 5.3.1, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
6 Conclusions
We have introduced a new text deduplication technique, LSHBloom, that offers significant runtime and space savings over the current state-of-the-art technique, MinHashLSH. We also conducted a thorough study of deduplication quality between our technique and other state-of-the-art deduplication algorithms commonly used in LLM training pipelines, using a custom benchmark of 25,000 scientific articles. We find that even with extensive parameter tuning for conventional MinHashLSH, LSHBloom is extremely competitive in terms of its deduplication quality, with an F1 score within 10% of the best achieved by MinHashLSH, while using far less storage and runtime. Using the peS2o dataset, we find that LSHBloom can feasibly scale to deduplicating billions of documents due to its low space requirements (a 180$\times$ space reduction from MinHashLSH) and relatively low wall clock time (3$\times$ speedup over MinHashLSH). In contrast, alternatives demonstrate either poor deduplication quality, excessive runtime, or prohibitive space requirements for extreme-scale datasets of this size. In the future, we aim to further reduce the wall clock time of LSHBloom by employing parallelization over subsets of text datasets when inserting them into our index—an approach that we expect to enable significant acceleration.

### v2

@@
 Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
-Figure 1: Breakdown of wall clock time on 5% of peS2o for a conventional LSH algorithm and our LSHBloom method
+Figure 1: Breakdown of wall clock time on 5% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
 2.4 Bloom Filter
@@
 We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix.
-By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for hash collisions.
+By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates.
 The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of Bender et al.
@@
 Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm.
-It is desirable to hash this list of integers onto a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions.
-To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function, with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values Carter & Wegman (1977).
-Then, we can hash a vector $\bar{x}$ of $r$ integers as follows:
+It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions.
+To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function (in our case, SHA1 3rd & Jones (2001)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values.
+Then, we can hash a vector $\bar{x}$ of $r$ integers as follows Carter & Wegman (1977):
 $h(\bar{x})=\left(\sum_{i=1}^{r}h_{i}(x_{i})\right)\mathrm{mod}\ {N}$
@@
 However, the false positive probability is configurable and can be set to be arbitrarily low (in our experiments, we set the effective false positive rate as low as 1e-5).
+4.4 Error Rate
+In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate.
+This is also the case in LSHBloom.
+However, we make use of Bloom filters, which incur additional false positive overhead.
+Here, we analyze the error rate of LSHBloom.
+A false positive is a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom, and a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
+[…]
+The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows.
+Leskovec et al.
+(2014) determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
+$FP=\int_{0}^{T}1-(1-t^{r})^{b}dt$ (1)
+$FN=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$ (2)
+Zhu et al. (2016) demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
+A Bloom filter will not yield false negatives, so the overall false negative rate is unaffected in LSHBloom.
+If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result.
+If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$.
+Therefore, the overall false positive rate for LSHBloom is given by:
+$FP_{total}=FP+(1-FP)*p_{effective}$
+Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible.
 4.5 Advantages
@@
 (2022).
-[…]
 As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing.
@@
 While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration.
-The additional false positive probability scales with the number of documents we intend to insert and scales inversely with the size of each filter.
+This is another advantage of our technique: we can observe massive gains in space while only introducing a single additional hyperparameter, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
+The additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter.
 Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter.
@@
 […]
+Figure 2: Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given $N$=24,956 documents and $p=$1e-5
 It is possible that one could properly tune these deduplication methods further using incremental changes in their parameter settings.
@@
 The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts.
-Figures 2(a) and 2(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset.
+Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset.
 As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate.
@@
 N-gram methods such as DOLMA-ngram and CCNet do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques.
-Figures 3(a) and 3 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method.
+Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method.
 While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus.
@@
 Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance.
-Figures 5(a) and 5(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively.
+Figures 6(a) and 6(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively.
 MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime.

### v3

@@
-
-Abstract
-Deduplication is a major focus for assembling and curating training datasets for large language models (LLM) – detecting and eliminating additional instances of the same content – in large collections of technical documents.
-Unrestrained, duplicates in the training dataset increase training costs and lead to undesirable properties such as memorization in trained models or cheating on evaluation.
-Contemporary approaches to document-level deduplication are often extremely expensive in both runtime and memory.
-We propose LSHBloom, an extension to MinhashLSH, which replaces the expensive LSHIndex with lightweight Bloom filters.
-LSHBloom demonstrates the same deduplication performance as MinhashLSH with only a marginal increase in false positives (as low as 1e-5 in our experiments); demonstrates competitive runtime (270% faster than MinhashLSH on peS2o); and, crucially, uses just 0.6% of the disk space required by MinhashLSH to deduplicate peS2o.
-We demonstrate that this space advantage scales with increased dataset size—at the extreme scale of several billion documents, LSHBloom promises a 250% speedup and a 54$\times$ space advantage over traditional MinHashLSH scaling deduplication of text datasets to many billions of documents.
-marginparsep has been altered.
-topmargin has been altered.
-marginparwidth has been altered.
-marginparpush has been altered.
-The page layout violates the ICML style.
-Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you.
-We’re not able to reliably undo arbitrary changes to the style.
-Please remove the offending package(s), or layout-changing commands and try again.
-LSHBloom: Memory-efficient, Extreme-scale Document Deduplication
-Anonymous Authors1
-footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
-Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
-1 Introduction
-Scholarly publications and related literature represent a tremendous body of often carefully encoded human knowledge.
-Google Scholar is estimated to index some 400 million documents Gusenbauer (2019), suggesting that the total number of unique documents worldwide numbers at least in the hundreds of millions.
-And a growing number of digital repositories mean that researchers can today access millions of such documents with ease, and with effort, tens of millions or more.
-Developments in large language models (LLMs) suggest that such corpora can be used to develop powerful agents able to assist with technical work in a variety of fields Zhang et al.
-(2024).
-However, seizing this opportunity at the largest scale requires addressing challenging deduplication problems.
-Any large corpus of technical documents assembled from diverse locations will include multiple copies of many documents: not only identical copies but also copies that vary, subtly or otherwise, for reasons such as different publication formats, the use of different parsers, copyediting, and versioning.
-As is well known, duplication in training datasets can lead to a host of undesirable behaviors, from excessive training times to predictive biases and training data memorization Lee et al.
-(2021) in trained models.
-Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets Chandran et al.
-(2024), illegitimately increasing perceived model performance by 8–11% Dekoninck et al.
-(2024).
-These considerations motivate the problem that we tackle in this paper, namely the accurate and efficient deduplication of large numbers (10s to 100s of millions) of scholarly and technical documents.
-We approach this problem from two perspectives.
-First, we demonstrate via careful evaluation that existing state-of-the-art approaches to the deduplication of technical documents suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs.
-Due to the frequent loss of metadata and variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method Broder et al.
-(1998).
-The infeasibility of pairwise comparisons in large datasets motivated the development of the MinHashLSH method Leskovec et al.
-(2014), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents.
-However, the associated MinHashLSH index can grow extremely large.
-For example, applying MinHashLSH to the peS2o dataset Soldaini & Lo (2023) of just 8M full-text and 31M abstracts takes more than 14 hours on a single 32-core node of the Polaris supercomputer and requires more than 300GB of disk space (see Section 5.3.1).
-Second, we introduce a new method, LSHBloom, that improves significantly on the state of the art in terms of the computational costs required to achieve a desired level of deduplication performance.
-The key insight here is that the process that MinHashLSH employs to find matches between MinHash signatures can be approximated, with a modest and controllable loss of accuracy, by using Bloom filters.
-This use of Bloom filters enables LSHBloom to run almost three times faster than MinHashLSH by interfacing with inexpensive bit arrays rather than inserting and querying against the comparatively slow tree data structures that comprise the traditional LSHIndex.
-Most importantly, this use of Bloom filters means that LSHBloom requires just 0.6% of the disk space consumed by the traditional MinHashLSH index to deduplicate peS2o Soldaini & Lo (2023), a dataset of 39 million documents.
-We demonstrate that this space advantage remains as the number of documents increases and makes it feasible for LSHBloom to process several billion documents with space savings of 54$\times$ relative to MinHashLSH without sacrificing deduplication quality.
-2 Background
+LSHBloom: Internet-Scale Text Deduplication
+Abstract.
+[…]
+PVLDB Reference Format:
+PVLDB, 19(1): XXX-XXX, 2025.
+doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
+doi:$X$.XX/XXX.XX
+1. Introduction
+Modern large language model (LLM) training pipelines require the ingestion of many millions of documents to enable internet-scale language pre-training.
+These documents are obtained through web scraping, deals with academic institutions, and relationships with publishers.
+The sheer scale of this data ingestion process presents a substantial data management and processing burden when assembling large-scale text databases for continuous LLM pre-training.
+LLM performance has also been shown to be sensitive to data quality (alignment of data with downstream tasks) (tan202415pintstechnicalreportpretraining) as well as data redundancy–excessive redundancy can cause performance scaling below the predicted rate of established scaling laws (chen-etal-2025-revisiting).
+Moreover, per-epoch training costs scale in the number of training data points.
+As such, it is highly desirable to remove outright uninformative data points from internet-scale text datasets–to alleviate the data management burden, to reduce the costs of pre-training, and to enhance LLM performance.
+In this work, we focus on reducing data redundancy in text.
+Duplicate text data–that is, instances of the same content or content that only subtly differ from one another due to differences in publication format, copyediting, and versioning–are ubiquitous in internet-scale text datasets and are well known to lead to a host of undesirable outcomes, from excessive training times to the introduction of predictive biases and training data memorization (lee2021deduplicating), in trained models.
+Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets (Chandran2024PrivateBT), illegitimately inflating perceived model performance by 8–11% (Dekoninck2024EvadingDC).
+These considerations motivate us to address the accurate and efficient deduplication of large numbers (100s of millions) of documents when preprocessing internet-scale text datasets.
+We demonstrate via careful evaluation that existing state-of-the-art approaches to text deduplication suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs.
+Due to the frequent loss of metadata and subtle variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method (broder1998min).
+The infeasibility of pairwise comparisons in large datasets motivated the development of MinHashLSH (leskovec2014finding), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents.
+MinHashLSH is by far the most popular method for text deduplication in LLM data pipelines; however, the associated MinHashLSH index can grow extremely large.
+For example, applying MinHashLSH to the peS2o dataset (peS2o) of just 39M documents takes more than 35 hours on a single 32-core node of the Polaris supercomputer and requires more than 200 GB of disk space (see subsubsection 5.4.1).
+[…]
+MinHashLSH is the most popular and performant state-of-the-art text deduplication technique available today; LSHBloom can be employed as a drop-in replacement for MinHashLSH–preserving this superior performance while enabling significant gains in scaling capacity as we increase our training dataset sizes.
+We know from historical data that LLM training dataset sizes will almost quadruple each year (epoch2024datasetsizetrend).
+LSHBloom presents an opportunity to create robust data processing infrastructure for LLM pre-training that is effective, efficient, and scalable well into the future.
+In summary, the contributions of our work are the following:
+  1. (1)
+    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data
+  2. (2)
+    We present the first study on the effects of tuning the hyperparameters for each of these algorithms, and present the best parameter settings for each
+  3. (3)
+    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime
+2. Background
 We first provide background information on the deduplication problem and describe the current state-of-the-art for large-scale text deduplication pipelines.
-2.1 Defining the Deduplication Problem
+2.1. Defining the Deduplication Problem
 Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus.
@@
 High computational performance is critical due to the extreme size of modern LLM datasets.
-For example, peS2o Soldaini & Lo (2023) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature Gusenbauer (2019), DOLMA Soldaini et al.
-(2024) contains roughly 5B documents, and RedPajama Computer (2023) has just over 100B documents—scales at which naive pairwise comparisons become infeasible.
+For example, peS2o (peS2o) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature (gusenbauer2019google), DOLMA (soldaini_dolma_2024) contains roughly 5B documents, and RedPajama (together2023redpajama) has just over 100B documents—scales at which naive pairwise comparisons become infeasible.
 Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
@@
 For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions.
-When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and various format-specific text parsing pipelines tend to induce small differences even when applied to the same underlying document.
+When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will induce small differences even when applied to the same underlying document.
 Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language.
 PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
-For these reasons, many deduplication techniques choose to take a “fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs.
-One way to do this is to alter one’s choice of similarity metric to encapsulate a broader semantic notion of similarity.
-This is common in so-called semantic deduplication techniques such as SemDeDup Abbas et al.
-(2023), which employ embeddings from pre-trained language models to identify semantically similar text data.
+For these reasons, many contemporary deduplication techniques take a ”fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs.
+One way to do this is to alter one’s choice of similarity metric to encapsulate a broader notion of similarity.
+This is common in so-called semantic deduplication techniques such as SemDeDup (abbas2023semdedup), which employ embeddings from pre-trained language models to identify semantically similar text data.
 Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents.
 It is more efficient to instead establish a similarity metric based on the raw text content of each document.
-2.2 Jaccard Similarity and MinHash Approximation
-Popular techniques such as those used to build the DOLMA Soldaini et al.
-(2024) and DataComp-LM Li et al.
-(2024) datasets measure n-gram overlap between documents to identify duplicates.
+2.2. Jaccard Similarity and MinHash Approximation
+Popular techniques such as those used to build the DOLMA (soldaini_dolma_2024) and DataComp-LM (li2024datacomp) datasets measure n-gram overlap between documents to identify duplicates.
 Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
-A text document can be viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
+In this paradigm, a text document is viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
 However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets.
 For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming.
-The MinHashing procedure Broder et al.
-(1998) provides a means of efficiently approximating the Jaccard similarity of two documents.
+The MinHashing procedure (broder1998min) provides a means of efficiently approximating the Jaccard similarity of two documents.
 Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures.
-The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets Broder et al.
-(1998); Leskovec et al.
-(2014).
-Therefore, we can repeatedly apply random permutations to the two sets’ elements and compare them for equality to obtain an estimate of the Jaccard similarity.
-Our estimate grows more accurate as we increase our sample size.
-2.3 MinHashLSH for Scalable MinHash
-The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed computed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document.
-However, the need to compare every pair of documents for similarity during deduplication still introduces a quadratic cost when comparing all documents in a corpus.
+The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (broder1998min; leskovec2014finding).
+Therefore, we can repeatedly apply the MinHashing procedure to any two documents and compare the resulting hashes for equality to obtain an estimate of their Jaccard similarity.
+This estimate grows more accurate as we increase our sample size.
+2.3. MinHashLSH for Scalable MinHash
+The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document.
+However, the need to compare every pair of documents for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
 Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs.
-MinHashLSH Leskovec et al.
-(2014) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents.
+MinHashLSH (leskovec2014finding) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents.
 Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document.
@@
 With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$.
-This enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity.
+This procedure enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity.
 This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
 One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication.
-However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 23 TB for 5B documents.
-Additionally, inserting and querying these tree-structured indices constitutes a significant portion (>80%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1).
+However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 277 TB for 5B documents.
+Additionally, inserting and querying these tree-structured indices constitutes a significant portion (¿85%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1).
 Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
-Figure 1: Breakdown of wall clock time on 5% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
-2.4 Bloom Filter
-The Bloom filter Bloom (1970) is a space-efficient probabilistic data structure used to test whether an element is a member of a set.
+Figure 1. Breakdown of wall clock time on 10% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
+2.4. Bloom Filter
+The Bloom filter (bloom1970space) is a space-efficient probabilistic data structure used to test whether an element is a member of a set.
 It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error.
@@
 If any of those $k$ bits is 0, the element is not contained in the set.
-Bloom filters have been used extensively in distributed computing for a variety of purposes Broder & Mitzenmacher (2004); Chervenak et al. (2008); Tarkoma et al. (2011).
-3 Related Work
+Bloom filters have been used extensively in distributed computing for a variety of purposes (broder2004network; chervenak2008globus; tarkoma2011theory).
+3. Related Work
 Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs.
 LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data.
-Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 Touvron et al.
-(2023b), Claude, Gemini Gemini Team (2024a; b), and GPT versions 3 and later OpenAI (2024).
-Others, such as Palm2 Anil et al.
-(2023) and Gemma Gemma Team (2024); Gemini Team (2024b), provide only limited details.
-Others, including several leading models, provide extensive details about how data for their model was dededuplicated including Olmo’s Dolma dataset Soldaini et al.
-(2024) as well as its successor DataComp-LM Li et al.
-(2024), Llama 1 Touvron et al.
-(2023a), Llama 3 Dubey et al.
-(2024), and OPT Zhang et al.
-(2022).
-For example, Llama 3 Dubey et al.
-(2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data.
-Phi uses an approach described by Gunasekar et al.
-(2023) and elaborated on in their Phi 2.0 report Javaheripi (2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation.
+Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 (touvron_llama_2023-1), Claude, Gemini (gemini_team_gemini_2024; gemini_team_gemini_2024-1), and GPT versions 3 and later (openai_gpt-4_2024).
+Others, such as Palm2 (anil2023palm) and Gemma (gemma_team_gemma_2024; gemini_team_gemini_2024-1), provide only limited details.
+Others, including several leading models, provide extensive details about how data for their model was deduplicated including Olmo’s Dolma dataset (soldaini_dolma_2024) as well as its successor DataComp-LM (li2024datacomp), Llama 1 (touvron_llama_2023), Llama 3 (dubey_llama_2024), and OPT (zhang_opt_2022).
+For example, Llama 3 (dubey_llama_2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data.
+Phi uses an approach described by gunasekar_textbooks_2023 and elaborated on in their Phi 2.0 report (javaheripi_surprising_2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation.
+C4 (t5_c4_raffel2020) applies only heuristic filtering for data quality (e.g., removing lines without punctuation or fewer than three words).
+RedPajama (together2023redpajama) uses both exact matching via Bloom filters and fuzzy deduplication via MinHashLSH.
+Datasets such as the Pile (gao2020pile800gbdatasetdiverse) and more recent releases such as FineWeb2 (penedo2025fineweb2pipelinescale) apply MinHashLSH for fuzzy deduplication.
 The most common methods used across these model sources include metadata-based approaches, MinHashLSH, CCNet, and n-gram-based methods.
@@
 ACM, different GitHub repositories); 2) sections/paragraphs of text are copied, as in extended journal versions of computer science papers; or 3) metadata is not curated (e.g., an ArXiV entry may not be updated to reflect a subsequent publication).
-CCNet Wenzek et al.
-(2020) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content.
+CCNet (wenzek2020ccnet) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content.
 CCNet performs dataset sharding, deduplication, language identification, and quality filtering.
@@
 CCNet can optionally compare to a single document, a subset of documents, or all documents.
-Dolma Soldaini et al.
-(2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma).
+Dolma (soldaini_dolma_2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma).
 An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter.
@@
 By default, $T$ = 1.0 in their configuration.
-DataComp-LM Li et al.
-(2024) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter.
-They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer Ye et al.
-(2023).
+DataComp-LM (li2024datacomp) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter.
+They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer (ye2023uniseg).
 For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams.
@@
 They employ the same threshold parameters for paragraph-level and document-level deduplication.
-4 The LSHBloom Method
-4.1 Overview
+A table comparing various methods for data deduplication in contemporary LLM pre-training pipelines can be found in Appendix A.
+4. The LSHBloom Method
+4.1. Overview
 We propose a memory-efficient alternative to MinHashLSH for document deduplication.
@@
 By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates.
-The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of Bender et al.
-(2022).
-4.2 Insertion
+The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of bender2022optimal.
+4.2. Insertion
 Given a Jaccard similarity threshold $T\in$ [0, 1], MinHashLSH traditionally groups signatures into a series of $b$ bands of size $r$ such that an exact match in one band for two columns of the MinHash signature matrix denotes a duplicate document.
@@
 It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions.
-To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function (in our case, SHA1 3rd & Jones (2001)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values.
-Then, we can hash a vector $\bar{x}$ of $r$ integers as follows Carter & Wegman (1977):
+To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (sha1)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values.
+Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (carter1977universal):
 $h(\bar{x})=\left(\sum_{i=1}^{r}h_{i}(x_{i})\right)\mathrm{mod}\ {N}$
-This approach enables us to map each band to a single integer value with known collision probability.
+This approach enables us to map each band of MinHash signatures to a single integer value with known collision probability ($1/N$).
 To insert a document into the index, we simply insert each of the $b$ final integer values into each of the $b$ Bloom filters corresponding to the $b$ bands in the signature matrix.
-4.3 Querying
+4.3. Querying
 Querying a document against the index is a similar process to insertion.
-We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in Section 4.2.
+We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in subsection 4.2.
 If there is a collision for any band of signatures, we mark that document as a duplicate.
 This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters.
-However, the false positive probability is configurable and can be set to be arbitrarily low (in our experiments, we set the effective false positive rate as low as 1e-5).
-4.4 Error Rate
+However, the false positive probability is configurable and can be set to be arbitrarily low.
+4.4. Error Rate
 In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate.
@@
 A false positive is a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom, and a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
-[…]
+Given a Minhash signature matrix with $b$ bands with $r$ rows each corresponding to $b$ Bloom filters with their individual false positive rates set to $p$, the net false positive overhead of using Bloom filters will be $p\textsubscript{effective}=1-(1-p)^{b}$.
+This is because a single collision in any band constitutes a duplicate, and false positive rates are independent for each Bloom filter.
+Therefore, the probability that no Bloom filter reports a false positive is $(1-p)^{b}$, and the overall false positive rate overhead introduced by using Bloom filters is the complement of this value.
 The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows.
-Leskovec et al.
-(2014) determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
-$FP=\int_{0}^{T}1-(1-t^{r})^{b}dt$ (1)
-$FN=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$ (2)
-Zhu et al. (2016) demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
-A Bloom filter will not yield false negatives, so the overall false negative rate is unaffected in LSHBloom.
+leskovec2014finding determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
+(1) $FP_{lsh}=\int_{0}^{T}1-(1-t^{r})^{b}dt$
+(2) $FN_{lsh}=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$
+zhu_lshensemble demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
+A Bloom filter will not yield false negatives, but may yield false positives.
 If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result.
 If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$.
+Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.2) for insertion into the Bloom filter.
 Therefore, the overall false positive rate for LSHBloom is given by:
-$FP_{total}=FP+(1-FP)*p_{effective}$
+(3) $FP_{bloom}=FP_{lsh}+(1-FP_{lsh})*(p_{effective}+b/N)$
 Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible.
-4.5 Advantages
+For 32-bit hashvalues (as is the default in packages such as datasketch) $N=2^{32}$, so the probability of collision for bands prior to insertion into the Bloom filter is also quite small.
+Additionally, because the size of the Bloom filter index scales with the number of items inserted and is not influenced by the size of the integers inserted into it, one can increase the size of the hashvalues (to say, a 64 or 128-bit integer) without increasing the size of the Bloom filter index.
+This is not true of the traditional MinHashLSH index which must store these hashvalues for comparison.
+Meanwhile, for each false negative that MinHashLSH produces, $p_{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the bloom filter index and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion.
+That is,
+(4) $FN_{bloom}=(1-(p_{effective}+b/N))*FN_{lsh}$
+4.5. Advantages
 We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams.
-Since this representation stems from the k-shingling of documents and many practical scenarios may admit an extremely large universe of such k-shingles (e.g., if we allow our set of documents to include a variety of encodings, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the range of MinHash hash values or increasing the accuracy of MinHash estimates of the Jaccard similarity.
-This can be done by increasing the number of permutations applied to prevent as much information loss as possible via compression of the text representation.
+Since many practical scenarios may admit an extremely large universe of such n-grams (e.g., if we allow our set of documents to include a variety of encodings, languages, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the size of MinHash hashvalues (to 64 or 128-bits or more) or increasing the accuracy of MinHash estimates of the Jaccard similarity by increasing the number of MinHash permutations.
 A scalable deduplication algorithm should be amenable to such considerations.
 LSHBloom allows for the space-efficient representation of extremely large deduplication indices.
-This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with the same number of bits.
-If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits Bender et al.
-(2022).
+This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with a fixed number of bits.
+This means that increasing the size of hashvalues does not change the index size.
+Furthermore, using more MinHash permutations only increases the index size linearly with the number of bands that are required for the MinHash signature matrix–which is far fewer than the number of permutations.
+If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (bender2022optimal).
 As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing.
-For these parameter settings, MinHashLSH creates nine bands for the MinHash signature matrix, meaning that we require nine Bloom filters for our index.
-For these settings, setting a conservative $p$effective = 1e-10 for a dataset of 10B documents would only require 590 GB of space.
-Meanwhile, traditional MinHashLSH may require 46 TB to store the equivalent LSHIndex—exemplifying the scalability of our technique.
+For these parameter settings, MinHashLSH creates nine bands (using the method described in (zhu_lshensemble) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index.
+For these settings, setting a conservative ppeffective = 1e-10 for a dataset of 10B documents would only require 590 GB of space.
+Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex—exemplifying the scalability of our technique.
 While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration.
@@
 However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.
-5 Evaluation
+5. Evaluation
 Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance.
 As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets.
-For correctness, we compare LSHBloom against state-of-the-art deduplication methods when applied to a dataset of 25K PDFs containing labeled duplicates.
-Without correctness, the system may either under-identify duplicates, resulting in decreased downstream inference quality; or over-identify duplicates, unnecessarily decreasing the volume of the training dataset.
-For performance, we consider the storage utilization and runtime of each method on a larger collection of PDFs, peS2o Soldaini & Lo (2023).
-Without performance, it may not be possible to process large datasets on available hardware resources.
-We conduct a comprehensive analysis of various deduplication methods.
-Specifically, as each method includes a set of parameters whose values affect both their correctness and resource usage, we evaluate each using a range of parameter settings.
-Understanding these relationships is crucial to making the best use of available resources.
+To assess correctness, we tune LSHBloom and other state-of-the-art deduplication methods on a balanced dataset of 25K PDFs containing labeled duplicates.
+We then use the best parameter settings for each algorithm and compare them on a larger labelled dataset of 50K PDFs.
+Without sufficient correctness, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary training overhead; or over-identifying duplicates, which unnecessarily reduces the training dataset volume.
+For performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (peS2o).
+Without sufficient performance, it may not be possible to process large datasets on available hardware resources.
 In the following subsections, we describe the details of our evaluation beginning with the evaluation methodology, then proceeding to correctness, performance, and finally the tradeoffs between performance and correctness.
-5.1 Experimental Methodology
+5.1. Experimental Methodology
 In this subsection, we describe our experimental methodology including testbed hardware, state-of-the-art methods for comparison, our evaluation metrics, and datasets.
-5.1.1 Testbed Hardware
+5.1.1. Testbed Hardware
 We ran our experiments on the Polaris supercomputer, a 560-node HPE Apollo 6500 Gen 10+-based system.
 Each node is equipped with a 32-core AMD Zen 3 2.8 GHz processor, with 512 GB of DDR4 RAM, and 3.2 TB of local SSD: representative of nodes found in large-scale high-performance computing facilities used for deduplication tasks.
-5.1.2 Baselines
+5.1.2. Baselines
 We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM.
 We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language.
-This also enabled us to use the same Bloom filter implementation for each technique.
+This also enabled us to normalize the Bloom filter implementation for those techniques which use them.
 Below, we describe any changes or configurations to the methods.
-MinhashLSH Broder (1997) (lsh): We configure the traditional MinhashLSH algorithm as described in Section 2.3.
+MinhashLSH (broder1997resemblance) (lsh): We configure the traditional MinhashLSH algorithm as described in subsection 2.3.
 DOLMA: We apply both standard dolma and dolma-ngram methods to document-level deduplication.
@@
 We parameterize Bloom filters for n-gram techniques using this estimate.
-5.1.3 Metrics
+We apply an analogous procedure to obtain estimated paragraph counts for Dolma and CCNet.
+5.1.3. Metrics
 Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents.
+Here, we consider a ”positive” label to mean that a document is a duplicate of another document in the corpus.
 We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance.
@@
 Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization.
-To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using the F1 score:
+To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using its precision (the proportion of duplicate labels that are true duplicates), its recall (the proportion of duplicates in the dataset that we correctly identify), and its F1 score (the harmonic mean of precision and recall):
 $F1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$
@@
 We measure wall clock time using /user/bin/time and assess each method’s disk space requirements using Python’s standard os.path.getsize method.
-5.1.4 Constructing the correctness evaluation dataset
-To permit the rigorous evaluation of de-duplication methods, we created an evaluation dataset By acquiring 24,000 scientific articles (4156 scraped from HTML and 20,844 parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata.
-We then duplicated 1000 of those documents for a total of 25,000.
-We then translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF Smith (2024), Nougat Blecher et al.
-(2023), Marker Paruchuri (2024), Grobid Lopez (2009), or Tesseract Smith (2007), each used with roughly the same frequency—for a total of 24,956 HTML or parsed PDF documents (as 44 failed to parse).
-Finally, we transformed 419 of these documents by removing a randomly selected subset of its tokens
-5.1.5 Choice of values for Tuning Parameters
-[…]
-Figure 2: Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given $N$=24,956 documents and $p=$1e-5
-It is possible that one could properly tune these deduplication methods further using incremental changes in their parameter settings.
+5.1.4. Constructing the correctness evaluation dataset
+To permit the rigorous evaluation of de-duplication methods, we created a pipeline for generating labeled datasets of duplicate documents.
+We acquire scientific articles (scraped from HTML and parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata.
+Text content from HTML is directly extracted.
+We translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF (PyMuPDF), Nougat (blecher2023nougat), or Tesseract (smith2007overview), each used with roughly the same frequency.
+We then can construct datasets where documents are duplicated, either by including multiple versions of the same document which were parsed using different text extraction tools and/or by including randomly truncated versions of that document.
+This reflects common forms of duplication that occur in real-world web scraping pipelines where the same document can be ingested via different text extraction tooling either due to differences in the origin of the document instance, the fact that the same document is often found in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors which can abruptly skip or truncate portions of the document’s text.
+We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm.
+We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances.
+5.1.5. Choice of values for Tuning Parameters
+We evaluate each technique over a range of evenly spaced parameter settings using a grid search.
+Specifically, for any threshold parameters such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values between 0.2 and 1.0 in steps of 0.2.
+We vary the number of permutations for MinHashLSH from 32 to 256 by powers of two.
+For n-gram techniques, we explore n-gram sizes 1, 2, 5, 7, 13, and 26.
+For MinHashLSH and LSHBloom, we additionally gathered results for $T=0.5$ and 48 random permutations as the surrounding parameter values demonstrated sharp jumps in the F1 score, and we wanted to explore this change in the score at a finer granularity.
+All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM, LSHBloom) are assigned a false positive rate $p$ of 1e-5.
+For LSHBloom, this equates to $p\textsubscript{effective}$ between 1e-5 and 8e-4, depending on the value of $T$ and the number of MinHash permutations.
+The value of $p_{effective}$ at various parameter settings for this benchmark is illustrated in Figure 2.
+Figure 2. Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given NN=24,956 documents and pp=1e-5
+It is possible that one could properly tune these deduplication methods further using more incremental changes in their parameter settings.
 However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance.
@@
 MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget.
-This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see Section 5.3.1).
+This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see subsubsection 5.4.1).
 (a) MinHashLSH
 (b) LSHBloom
-Figure 3: F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
-5.2 Deduplication Correctness
-5.2.1 LSH Methods
+Figure 3. F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
+5.2. Hyperparameter Tuning Results (24K Dataset)
+5.2.1. LSH Methods
 The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts.
-Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset.
-As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate.
-However, there is a clear saturation point at 128 permutation functions, after which we observe only marginal performance increases.
-This is particularly relevant for LSHBloom, for which increasing the number of permutation functions blindly can lead to slight degradations in performance—due to the introduction of more bands in the signature matrix by the LSH function, requiring more Bloom filters, and therefore leading to an increase in the false positive rate.
-This can be avoided, however, by jointly tuning the false positive rate and the number of permutation functions.
+Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 24K dataset.
+As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (broder1998min).
+Correspondingly, we use 256 MinHash permutations in our subsequent experiments.
 The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content.
-For our particular benchmark, a threshold of 0.6 yields the best performance, as more stringent thresholds lead to a diminishing F1 score.
+For our particular benchmark, a threshold of 0.5 yields the best performance; more stringent thresholds lead to a diminishing F1 score.
 (a) DOLMA
 (b) DataComp-LM
-Figure 4: F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
-5.2.2 Paragraph Level and N-Gram Methods
+Figure 4. F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
+5.2.2. Paragraph Level and N-Gram Methods
 Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score.
 Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance.
-N-gram methods such as DOLMA-ngram and CCNet do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques.
+The best F1 for paragraph-level methods is achieved by setting a low threshold of 0.2.
+N-gram methods such as DCLM do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques.
+Interestingly, DOLMA and DOLMA-ngram perform very similarly.
 Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method.
 While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus.
-This is similar to the notion of evaluating the set overlap between documents but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error.
-Moreover, both paragraph and n-gram methods demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions.
+This is similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error.
+Moreover, paragraph-level techniques demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions.
 Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates.
-In contrast, n-gram methods yield many false negatives at high thresholds and yield many false positives at low thresholds.
-Figure 5: F1 score vs. threshold for paragraph-level techniques.
-5.3 Deduplication Resource Usage
+Table 1 shows the parameter settings for each technique that we use for the remainder of our experiments.
+Figure 5. F1 score vs. threshold for paragraph-level techniques.
+Table 1. Best Settings for Each Deduplication Technique
+Technique N-gram Size Threshold
+MinHashLSH 1 0.5
+LSHBloom 1 0.5
+Dolma-Ngram 5 0.2
+DCLM 5 0.2
+Dolma - 0.2
+CCNet - 0.2
+5.3. Deduplication Correctness (50K Dataset)
+(a) Precision
+(b) Recall
+(c) F1
+Figure 6. Correctness metrics for deduplication techniques for various levels of duplication.
+5.3.1. Deduplication Quality
+Here we evaluate the correctness of each technique using metrics such as their precision, recall, and F1 score.
+We do so by constructing datasets of 50,000 documents each with varying levels of duplication between 10% and 90%.
+We construct these datasets using the procedure described in subsubsection 5.1.4.
+Figure 6 illustrates the precision, recall, and F1 score of each technique as we vary the proportion of duplicate documents in our dataset.
+Notably, in terms of F1 score (see 6(c)), both MinHashLSH and LSHBloom outperform all other methods—only being slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority ($60\%$) of the dataset consists of duplicates, which is highly unlikely in real-world scenarios.
+MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match the performance of MinHashLSH while affording considerable gains in scalability (to be shown below in subsection 5.4).
+Looking more closely, 6(a) illustrates that MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset.
+The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision often tracking closely with that of the MinHash-based techniques.
+However, Dolma’s dismal recall scores (<0.4<0.4) suggest that this is likely the result of under-identifying duplicate instances rather than being a reflection of greater capacity for discernment.
+DCLM, Dolma-Ngram, and CCNet tend to underperform the other techniques in terms of precision.
+Note that there is a natural tendency for precision scores to increase as the proportion of duplicates in the dataset increases.
+This is due to a greater number of true positives being present in the dataset.
+In terms of recall, 6(b) demonstrates that DCLM and Dolma-Ngram tend to perform the best—though notably their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare.
+This indicates that, while these techniques identify most duplicates, they struggle to differentiate between duplicate and non-duplicate entities.
+Meanwhile, MinHashLSH and LSHBloom provide virtually identical recall scores to one another between 0.7 and 0.9 while simultaneously demonstrating strong precision.
+Paragraph-level techniques demonstrate lower recall than all other techniques by a wide margin.
+Our experimental results indicate that LSHBloom achieves deduplication performance comparable to the state-of-the-art—indeed, LSHBloom achieves essentially identical performance to MinHashLSH, one of the most popular choices for text deduplication, with only a marginal degradation in precision.
+5.3.2. Performance vs. Quality Tradeoffs
+Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance.
+We visualize data from our evaluation of the balanced dataset of 50,000 documents.
+Figures 7(a) and 7(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively.
+MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime.
+LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see subsection 5.4).
+LSHBloom also displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index.
+As shown below in subsection 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
+Our results indicate that LSHBloom intelligently discerns between duplicate and non-duplicate documents (as demonstrated by its high precision scores) and can adeptly identify the majority of duplicate documents (as demonstrated by its strong recall scores).
+Our results also illustrate that the main difference in performance between MinHashLSH and LSHBloom comes in the form of slightly reduced precision for the latter.
+Intuitively, this makes sense as LSHBloom yields slightly more false positive predictions due to its use of the Bloom filter.
+Overall, however, both yield similar F1 scores.
+This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication use-cases.
+5.4. Deduplication Resource Usage at Scale
 (a) Quality vs. Wall Clock Time
 (b) Quality vs. Disk
-Figure 6: Pareto plots for F1 score vs. resource usage.
+Figure 7. Pareto plots for F1 score vs. resource usage.
 As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents.
-Here we work with the peS2o dataset Soldaini & Lo (2023) of roughly 8M full-text and 31M abstracts.
+Here we work with the peS2o dataset (peS2o) of roughly 39M documents.
 We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter).
 We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost.
-DataComp-LM, for example, takes one hour just to deduplicate 25,000 examples in our previous benchmark in Section 5.1.4.
-Even processing just 10% of peS2o using this technique would require 160 hours by extrapolation.
+DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in subsection 5.3.
+Even processing just 10% of peS2o using this technique would require over 200 hours by extrapolation.
 We discuss the results for other methods below.
-Figure 7: Disk size of index for different methods, as measured for peS2o subsets of increasing size.
-Figure 8: Wall clock times for different methods, as measured for peS2o subsets of increasing size.
-5.3.1 Scaling of Resource Usage
+Figure 8. Disk size of index for different methods, as measured for peS2o subsets of increasing size.
+Figure 9. Wall clock times for different methods, as measured for peS2o subsets of increasing size.
+5.4.1. Scaling of Resource Usage
 In this section, we study how resource usage grows with the number of documents.
-In these experiments, we set any threshold parameters to 0.8, the number of random permutations for our LSH techniques to 128, and $p$ = 1e-5 for all Bloom filters.
-Ideally, we would have informed these parameter settings based on the optimal settings from Section 5.1.4; however, scaling experiments on peS2o regularly took over a day to complete, limiting our ability to tune parameters and necessitating heuristic choices for parameter settings as we evaluated deduplication quality in parallel.
-Figure 8 shows the wall clock time to deduplicate various subsets of peS2o with each method.
-Notably, MinHashLSH incurs the largest runtime while paragraph-level techniques such as DOLMA and CCNet are relatively fast.
-However, DOLMA and CCNet achieve this speed with significantly lower deduplication performance: see Section 5.1.4.
-Meanwhile, LSHBloom, which has comparable deduplication performance to MinHashLSH, is much faster.
-To deduplicate all of peS2o, LSHBloom takes just five hours, whereas MinHashLSH takes 14 hours—almost a 3$\times$ speedup.
-Figure 1 shows the breakdown of this time for MinHashLSH and LSHBloom; we can see that, as expected, the time to compute MinHashes is the same; however, the insertion and query latencies are significantly reduced in the Bloom filter-based index compared to the traditional LSHIndex.
-Figure 7 shows the disk usage for each method.
-We again see a similar trend.
-The paragraph-level techniques maintain only a single Bloom filter which leads to extremely low disk usage.
-Meanwhile, MinHashLSH’s index grows steadily in size as the number of documents increases.
-LSHBloom’s index also grows linearly in the number of documents but at a reduced rate.
-When processing all of peS2o, LSHBloom requires just 1 GB vs.
-the 180 GB taken by MinHashLSH—a 180$\times$ reduction.
-5.3.2 Estimated Resource Usage at Scale
-Figure 9: Predicted wall clock times vs. number of documents (in Billions) for different methods
+In these experiments, we use the same settings shown in Table 1.
+We set the effective false positive rate for all bloom filters to 1e-10.
 […]
-Table 1: Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
-Technique Bloom filter false Index disk
-positive overhead usage (GB)
-MinHashLSH 23,008.04
-LSHBloom 1e-5 160.51
-LSHBloom 1e-10 295.30
-LSHBloom 1e-15 425.35
-Table 2: Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
+[…]
+5.4.2. Estimated Resource Usage at Scale
+Figure 10. Predicted wall clock times vs. number of documents (in Billions) for different methods
+[…]
+[…]
+Table 2. Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
 Technique Bloom filter false Index disk
 positive overhead usage (TB)
-MinHashLSH 460.16
-LSHBloom 1e-5 3.21
-LSHBloom 1e-10 5.91
-LSHBloom 1e-15 8.51
-5.4 Performance vs. Quality Tradeoffs at Scale
-Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance.
-Figures 6(a) and 6(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively.
-MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime.
-Additionally, LSHBloom displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index.
-As shown in Section 5.3.1, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
-6 Conclusions
-We have introduced a new text deduplication technique, LSHBloom, that offers significant runtime and space savings over the current state-of-the-art technique, MinHashLSH.
-We also conducted a thorough study of deduplication quality between our technique and other state-of-the-art deduplication algorithms commonly used in LLM training pipelines, using a custom benchmark of 25,000 scientific articles.
-We find that even with extensive parameter tuning for conventional MinHashLSH, LSHBloom is extremely competitive in terms of its deduplication quality, with an F1 score within 10% of the best achieved by MinHashLSH, while using far less storage and runtime.
-Using the peS2o dataset, we find that LSHBloom can feasibly scale to deduplicating billions of documents due to its low space requirements (a 180$\times$ space reduction from MinHashLSH) and relatively low wall clock time (3$\times$ speedup over MinHashLSH).
-In contrast, alternatives demonstrate either poor deduplication quality, excessive runtime, or prohibitive space requirements for extreme-scale datasets of this size.
+MinHashLSH 277.68
+LSHBloom 1e-5 8.33
+LSHBloom 1e-8 12.11
+LSHBloom 1/5B 15.50
+Table 3. Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
+Technique Bloom filter false Index disk
+positive overhead usage (TB)
+MinHashLSH 555.35
+LSHBloom 1e-5 16.66
+LSHBloom 1e-8 24.21
+LSHBloom 1/100B 31.76
+6. Conclusions
+[…]
 In the future, we aim to further reduce the wall clock time of LSHBloom by employing parallelization over subsets of text datasets when inserting them into our index—an approach that we expect to enable significant acceleration.
+Additionally, though very useful, our synthetic benchmark has drawbacks—namely, in its biased and limited notion of what constitutes a proper duplicate.
+Ideally, future work will involve benchmark datasets that cover more diverse forms of duplication in proportion to their representation ”in the wild”—that is, in real-world LLM pre-training data pipelines.
+Appendix A Comparison of Data Deduplication Pipelines for LLM Pre-training
+Table 4. Comparative Analysis of Data Deduplication Pipelines for LLM Pre-training
+Dataset / Model Primary Data Sources Key Cleaning & Filtering Steps Dedup. Level Deduplication Technique(s) Key Parameters & Implementation Notes Citations
+C4 (for T5) Common Crawl (April 2019 snapshot) Heuristic filtering: remove lines without terminal punctuation, lines << 3 words, docs << 5 sentences, docs with ”Lorem ipsum”, and docs containing words from a blocklist.
+Language filtering for English (>>99% probability).
+None Heuristic Filtering Only No explicit deduplication algorithms were used.
+The focus was on rule-based filtering.
+They remove all but one instance of any three-sentence span that occurs more than once in the dataset.
+(t5_c4_raffel2020)
+The Pile (for GPT-Neo) 22 diverse sources including Common Crawl (Pile-CC), Books3, arXiv, GitHub, OpenWebText2, and PubMed Central.
+Source-specific cleaning (e.g., converting TeX to Markdown for arXiv, parsing top answers for StackExchange).
+The entire dataset was filtered for duplicates.
+Document MinHashLSH The OpenWebText2 component specifically used in-memory MinHashLSH via the datasketch library for near-duplicate detection.
+(gao2020pile800gbdatasetdiverse)
+GPT-3 A mix of Common Crawl (filtered), WebText2, Books1, Books2, and Wikipedia.
+Quality filtering was applied to the Common Crawl portion.
+Document MinHashLSH Spark’s MinHashLSH implementation with 10 hashes.
+(brown2020language)
+RedPajama / SlimPajama (LLaMA Replication) LLaMA sources: Common Crawl, C4, GitHub, Books, arXiv, StackExchange.
+Multi-stage pipeline: NFC Unicode normalization, filtering documents << 200 characters.
+Document Exact Hashing (Bloom Filter) & Fuzzy Deduplication (MinHashLSH) RedPajama-V2 uses a Bloom filter for efficient exact deduplication in addition to MinHashLSH.
+SlimPajama uses a distributed MinHashLSH implementation with 13-grams.
+(together2023redpajama; cerebras2023slimpajama)
+CCNet CommonCrawl Language identification (for splitting into monolingual datasets) using lightweight classifiers.
+LM-based quality filtering at the document level.
+Exact (SHA1 Hashing) Line-level Computes SHA1 hashes for each line of text and performs exact deduplication at the line-level using these hash values.
+(wenzek2020ccnet)
+Llama 3 Constructed ”from a variety of data sources containing knowledge until the end of 2023”.
+PII and safety filtering.
+Heuristic filtering: removed lines with repeated content, removed lines with ”dirty” words, filtered out documents with excessive numbers of outlier tokens compared to the rest of the training corpus.
+Document and Line-Level URL (metadata), MinHash, CCNet The paper mentions the use of URL deduplication, MinHash for document-level deduplication, and CCNet for line-level deduplication.
+(dubey_llama_2024)
+Dolma Collected from CommonCrawl, Github, Reddit, Wikipedia, among other sources.
+Remove data containing PII or toxic material.
+Not Specified Paragraph Deduplicates paragraphs via exact-matching (Bloom filter).
+Alternatively, deduplicates paragraphs by measuring their n-gram overlap with the rest of the corpus, also using a Bloom filter.
+Preprocess some web data using CCNet and metadata-based deduplication.
+(soldaini_dolma_2024)
+DataComp-LM (DCLM) Common Crawl (provides a candidate pool of 240T+ tokens).
+A benchmark for data curation strategies.
+The DCLM-Baseline dataset was created using model-based filtering (data quality classifiers) to select high-quality data.
+Paragraph and Document-level Baseline dataset is deduplicated using Bloom filters to measure n-gram overlap of a paragraph/document with the rest of the dataset.
+The DCLM-Baseline dataset of ~4T tokens is combined with StarCoder and ProofPile2.
+(li2024datacomp)
+FineWeb / FineWeb-Edu / FineWeb2 Common Crawl (96 snapshots for FineWeb, ~100 for FineWeb2).
+URL filtering (blocklist), language classification (fastText), and quality/repetition filters.
+FineWeb-Edu uses a classifier trained on Llama-3-70B annotations.
+Document Fuzzy Deduplication (MinHash) Extensive ablation studies.
+The final strategy involved deduplicating each Common Crawl snapshot *individually*, as global deduplication was found to discard high-quality data from older snapshots.
+(penedo2025fineweb2pipelinescale)
+Lee et al.
+(2021) C4, RealNews, LM1B, Wiki-40B-en N/A (focus was on the deduplication technique itself).
+Substring (Exact) Suffix Array Removed all but one occurrence of any 50-gram (sequence of 50 tokens) that appeared in multiple documents.
+This method targets boilerplate text.
+(lee2021deduplicating)
+D4 (SemDeDup) Common Crawl (CC-dedup) An initial pass of MinHash-based deduplication was applied before the main technique.
+Document (Semantic) Semantic Deduplication (Embeddings + K-Means) A multi-stage method that first applies semantic deduplication (SemDeDup) to remove dense clusters of semantically redundant information.
+(tirumala2023d4)
+LSHBloom peS2o (a dataset of 39 million technical documents) N/A (focus was on the deduplication technique itself).
+Document (Approximate) MinHashLSH with Bloom Filter Index An extension of MinHashLSH that replaces the large, memory-intensive LSH index with a series of lightweight Bloom filters (one for each band).
+$12\times$ faster, $18\times$ less space usage.

### v4

@@
 Abstract.
-[…]
+Modern internet-scale text databases require robust entity deduplication to maintain data integrity and reduce processing overhead.
+While Locality-Sensitive Hashing (LSH) is the established paradigm for near-duplicate detection, traditional LSHIndex structures based on trees or hashmaps suffer from large disk footprints and pointer-chasing latencies.
+We propose LSHBloom, a novel architectural extension to MinHashLSH that replaces traditional indexing structures with an array of independent probabilistic Bloom filters.
+By mapping LSH bands directly to contiguous bit vectors, optimizing for spatial locality, and co-designing our software with modern supercomputer hardware, LSHBloom achieves $12\times$ higher throughput than standard MinHashLSH on the 39M-document peS2o dataset.
+Our design reduces disk footprint by $18\times$ while maintaining near-identical recall and precision.
+We provide a formal analysis of the analytically boundable error rates and demonstrate, through extrapolation, that LSHBloom enables billion-scale record deduplication on commodity hardware.
 PVLDB Reference Format:
@@
 1. Introduction
-Modern large language model (LLM) training pipelines require the ingestion of many millions of documents to enable internet-scale language pre-training.
-These documents are obtained through web scraping, deals with academic institutions, and relationships with publishers.
-The sheer scale of this data ingestion process presents a substantial data management and processing burden when assembling large-scale text databases for continuous LLM pre-training.
-LLM performance has also been shown to be sensitive to data quality (alignment of data with downstream tasks) (tan202415pintstechnicalreportpretraining) as well as data redundancy–excessive redundancy can cause performance scaling below the predicted rate of established scaling laws (chen-etal-2025-revisiting).
-Moreover, per-epoch training costs scale in the number of training data points.
-As such, it is highly desirable to remove outright uninformative data points from internet-scale text datasets–to alleviate the data management burden, to reduce the costs of pre-training, and to enhance LLM performance.
-In this work, we focus on reducing data redundancy in text.
-Duplicate text data–that is, instances of the same content or content that only subtly differ from one another due to differences in publication format, copyediting, and versioning–are ubiquitous in internet-scale text datasets and are well known to lead to a host of undesirable outcomes, from excessive training times to the introduction of predictive biases and training data memorization (lee2021deduplicating), in trained models.
-Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets (Chandran2024PrivateBT), illegitimately inflating perceived model performance by 8–11% (Dekoninck2024EvadingDC).
-These considerations motivate us to address the accurate and efficient deduplication of large numbers (100s of millions) of documents when preprocessing internet-scale text datasets.
-We demonstrate via careful evaluation that existing state-of-the-art approaches to text deduplication suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs.
-Due to the frequent loss of metadata and subtle variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method (broder1998min).
-The infeasibility of pairwise comparisons in large datasets motivated the development of MinHashLSH (leskovec2014finding), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents.
-MinHashLSH is by far the most popular method for text deduplication in LLM data pipelines; however, the associated MinHashLSH index can grow extremely large.
-For example, applying MinHashLSH to the peS2o dataset (peS2o) of just 39M documents takes more than 35 hours on a single 32-core node of the Polaris supercomputer and requires more than 200 GB of disk space (see subsubsection 5.4.1).
+Modern text databases must handle numerous texts at the internet-scale, particularly with the advent of data-hungry systems like large language models (LLMs) that require access to many millions of documents.
+Redundant text records are a major source of inefficiency in these systems—recent studies have demonstrated that large web-scraped corpora such as CommonCrawl (Nagel, 2023) (roughly 300B webpages) are comprised of between 14% to 52% near-duplicate records (Fröbe et al., 2021).
+Many modern applications that rely on text databases benefit strongly from procedures that mitigate data redundancy including 1) internet-scale LLM pre-training where data redundancy has marked negative effects on language modeling ability (Tan and Wang, 2024; Lee et al., 2021) and significantly inflates training costs (Chen et al., 2025), 2) retrieval-augmented generation systems that utilize vector databases where storing $O(d)$ fewer documents results in $O(log(D))$ faster query times (Malkov and Yashunin, 2018) and mitigates the substantial GPU cost involved in embedding text for insertion (Gao et al., 2023), 3) full-text search systems where storing $O(d)$ fewer documents results in a linear decrease in storage costs and in query complexity (Zobel and Moffat, 2006; Ayyagiri et al., 2024), 4) database tasks such as topic modeling or clustering where $O(d)$ fewer documents translates into an $O(d)$ speedup (Blei et al., 2003; Grootendorst, 2022), and 5) lookup tasks, such as those in search engines, that benefit from operations that can group text data by near-duplicates to group or distill query results (e.g., near-duplicate detection, fuzzy grouping) (Charikar, 2002; Andoni et al., 2018; Broder, 1997).
+Traditional relational entity resolution techniques based on exact-matching or common field values (e.g., blocking) (Christen, 2012) fall short for text due to parsing noise (aberrations in raw text content introduced when parsing from PDF, XML, HTML, etc.
+into raw text) and the frequent absence or unreliability of metadata when obtaining documents from the web.
+Consequently, recent work has focused on the use of approximate similarity metrics such as n-gram overlap (Soldaini and others, 2024; Li et al., 2024) and the Jaccard similarity (Broder, 1997; Broder et al., 1998).
+The computational infeasibility of performing pairwise comparisons between documents in large text datasets motivated the development of MinHashLSH (Leskovec et al., 2014), which applies a locality-sensitive hashing (LSH) scheme to avoid comparing dissimilar documents.
+MinHashLSH is by far the most popular method for text deduplication in LLM data ingestion workflows (Brown et al., 2020; Rae et al., 2021; Olmo et al., 2025; Touvron et al., 2023; Dubey and others, 2024; Computer, 2023; Gao et al., 2020; Mireshghallah et al., 2023; Black et al., 2022; Biderman et al., 2023; Penedo et al., 2025, 2023; Li et al., 2023; Lozhkov et al., 2024); however, the associated MinHashLSH index can become slow and grow extremely large when applied at internet scale, presenting bottlenecks in throughput as well as in memory and disk resources.
+For example, applying MinHashLSH to the tiny peS2o dataset (Soldaini and Lo, 2023) containing 39M academic documents takes more than 35 core hours on the Polaris supercomputer and requires more than 200 GB of disk space (see Section 5.4.1).
+Overcoming these bottlenecks is paramount: language models are routinely retrained with newly acquired data (Dai et al., 2024; Parmar et al., 2024) and vector databases are continually updated with new content (Taipalus, 2024; Ma et al., 2023).
+Continuously updating these collections requires efficient data ingestion, cleaning, and deduplication workflows.
+Moreover, when constructing text datasets, upstream changes in data pipelines such as the mode of data ingestion, cleaning, or preprocessing may alter the representation of existing documents and therefore, necessitate rerunning deduplication workflows on large corpora to mitigate data redundancy.
+Beyond a certain size, the runtime and disk requirements imposed by existing state-of-the-art deduplication techniques, such as MinHashLSH, render these processes infeasible to run.
 […]
-MinHashLSH is the most popular and performant state-of-the-art text deduplication technique available today; LSHBloom can be employed as a drop-in replacement for MinHashLSH–preserving this superior performance while enabling significant gains in scaling capacity as we increase our training dataset sizes.
-We know from historical data that LLM training dataset sizes will almost quadruple each year (epoch2024datasetsizetrend).
-LSHBloom presents an opportunity to create robust data processing infrastructure for LLM pre-training that is effective, efficient, and scalable well into the future.
-In summary, the contributions of our work are the following:
+LSHBloom is designed as a drop-in replacement for the industry-standard MinHashLSH (Brown et al., 2020; Olmo et al., 2025; Dubey and others, 2024; Rae et al., 2021)—preserving deduplication fidelity while providing significant headroom in scaling capacity as the cardinality of text datasets increases.
+Historical analysis of frontier models indicates that LLM training dataset sizes almost quadruple each year (Rahman and Owen, 2024).
+Beyond the increasing availability of large-scale general-purpose web corpora (Gao et al., 2020; Penedo et al., 2023, 2025) (including CommonCrawl (Nagel, 2023), which adds 3-5B newly scraped web pages each month), coordinated national initiatives such as the Genesis Mission (Executive Office of the President, 2025) will integrate decades of scientific research into new, large-scale datasets for LLM training.
+This trend towards larger text datasets is further exacerbated by the recent strategic shift towards using high-volume synthetic data generation to continue scaling LLM training (Gunasekar et al., 2023; Javaheripi, 2023) despite limitations in the amount of available human-authored data.
+The ease with which one can generate vast quantities of synthetic text contributes to an ever-increasing data management burden.
+These observations motivate the development of LSHBloom, a robust data deduplication technique for large text datasets that is effective, efficient, and scalable well into the future.
+In summary, our contributions are:
   1. (1)
-    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data
+    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized by LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data.
   2. (2)
-    We present the first study on the effects of tuning the hyperparameters for each of these algorithms, and present the best parameter settings for each
+    We present an early study on the effects of tuning the hyperparameters for each of these algorithms, and identify the best parameter settings for each.
   3. (3)
-    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime
+    We carefully co-design the implementation with the modern supercomputer hardware to maximize throughput.
+  4. (4)
+    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime.
 2. Background
-We first provide background information on the deduplication problem and describe the current state-of-the-art for large-scale text deduplication pipelines.
-2.1. Defining the Deduplication Problem
-Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus.
-These methods can be applied at different resolutions (e.g., to documents, paragraphs, sentences, or even tokens) and scopes (e.g., across a corpus or within a single document).
-Deduplication methods may also differ in how they define duplication (e.g., exact matches, n-gram overlap, “fuzzy”).
-In evaluating a deduplication method for a particular application, we are concerned with its precision, recall, and computational performance.
-Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces the size of the training dataset and limits an LLM’s ability to learn the underlying distribution of the text data.
-Low recall, i.e., excessive false negatives, admits a great degree of duplicate content in our dataset, with negative effects on learning as well as potentially inducing degenerate behavior such as memorization.
-High computational performance is critical due to the extreme size of modern LLM datasets.
-For example, peS2o (peS2o) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature (gusenbauer2019google), DOLMA (soldaini_dolma_2024) contains roughly 5B documents, and RedPajama (together2023redpajama) has just over 100B documents—scales at which naive pairwise comparisons become infeasible.
-Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
-Here, we focus on methods for document-level deduplication.
-Suppose an operator $S(d_{i},d_{j}):D\times D\rightarrow[0,1]$ that takes as input two documents and outputs a similarity score between 0 and 1.
-We define our task as follows: Given a corpus of $N$ documents, $D_{N}=$ $d_{1}$, …, $d_{N}$}, and a similarity threshold, $T\in[0,1]$, identify every document $d$ such that $\exists d^{\prime}\in D_{N}$ such that $d^{\prime}\neq d$ and $S(d,d^{\prime})\geq T$.
-Just how to define duplication is a subtle question.
-For example, one could stipulate simply that two documents must be byte-for-byte identical: equivalent to setting $S$ to measure byte overlap and $T$ = 1.0 above.
-However, this definition may exclude many document pairs that we would like to consider duplicates due to great similarity in their token distributions or semantic content, as are encountered frequently in practical text ingestion pipelines.
+We first introduce the deduplication problem and describe the state-of-the-art for large-scale text deduplication pipelines.
+2.1. Defining the Deduplication Problem for Text
+In the context of database systems, text deduplication is a specialized form of Online Duplicate Elimination.
+Unlike batch-oriented entity resolution, which materializes all pairwise similarities within a static table (Christen, 2012), text deduplication in modern ingestion pipelines functions as a Streaming Approximate Membership Query (SAMQ).
+The objective is to determine, in real-time, if an incoming record represents a redundant entry relative to the existing corpus state.
+Formally, we define this task as a set-similarity membership problem.
+Suppose a similarity function $S(d_{i},d_{j}):D\times D\rightarrow[0,1]$ that takes as input two documents and outputs a similarity score between 0 and 1.
+We define our task as follows: Given a stream of $N$ documents, $D_{N}=$ $d_{1}$, …, $d_{N}$}, and a similarity threshold, $T\in[0,1]$, for each document, $d_{i}$, evaluate the binary decision function $\mathcal{F}(d_{i})$ which returns 1 if $\exists d_{j}\in D_{seen}:S(d_{i},d_{j})\geq T$ where $D_{seen}=\{d_{j}\mid j<i\}$ and 0 otherwise.
+This formulation prioritizes ingestion throughput and state compression as storing the full history, $D_{seen}$, is intractable at scale.
+The choice of similarity function, $S$, is crucial for maximizing join fidelity.
+For example, one could simply stipulate that two documents must be byte-for-byte identical.
+However, this definition may exclude many document pairs that are not exact matches, but have high overlap in their content, as are encountered frequently in practical text ingestion pipelines.
 For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions.
-When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will induce small differences even when applied to the same underlying document.
-Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language.
+Also, when web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will produce small differences in raw text content even when applied to the same underlying document.
 PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
-For these reasons, many contemporary deduplication techniques take a ”fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs.
-One way to do this is to alter one’s choice of similarity metric to encapsulate a broader notion of similarity.
-This is common in so-called semantic deduplication techniques such as SemDeDup (abbas2023semdedup), which employ embeddings from pre-trained language models to identify semantically similar text data.
-Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents.
-It is more efficient to instead establish a similarity metric based on the raw text content of each document.
-2.2. Jaccard Similarity and MinHash Approximation
-Popular techniques such as those used to build the DOLMA (soldaini_dolma_2024) and DataComp-LM (li2024datacomp) datasets measure n-gram overlap between documents to identify duplicates.
-Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
-In this paradigm, a text document is viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
-However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets.
-For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming.
-The MinHashing procedure (broder1998min) provides a means of efficiently approximating the Jaccard similarity of two documents.
-Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures.
-The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (broder1998min; leskovec2014finding).
+As a result, properly deduplicating text data requires adopting approximate similarity measures that are robust to minor aberrations.
+2.2. Approximate Similarity Metrics for Text
+One way to define an approximate similarity measure between documents is to measure the proportion of n-grams in that document that were previously seen in the data stream.
+Popular techniques such as those used to build the DOLMA (Soldaini and others, 2024) dataset and DataComp-LM (Li et al., 2024) baseline dataset employ this procedure to identify duplicates.
+However, these techniques can be sensitive to the relative frequency of n-grams within a document or dataset.
+Alternatively, some techniques, such as MinHashLSH, take a set-theoretic perspective on document similarity.
+In this set-theoretic paradigm, a text document is viewed as a set of n-grams, and the similarity between two such sets, $A$ and $B$, is defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
+However, computing the Jaccard similarity requires evaluating set membership for each item present in the union of both sets.
+For extremely high-dimensional objects such as documents, computing the Jaccard similarity explicitly can be extremely expensive.
+The MinHashing procedure (Broder et al., 1998) provides a means of efficiently approximating the Jaccard similarity of two documents.
+Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures—lossy, fixed-length representations of the underlying high-dimensional objects.
+The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (Broder et al., 1998; Leskovec et al., 2014).
 Therefore, we can repeatedly apply the MinHashing procedure to any two documents and compare the resulting hashes for equality to obtain an estimate of their Jaccard similarity.
-This estimate grows more accurate as we increase our sample size.
-2.3. MinHashLSH for Scalable MinHash
+This estimate grows more accurate as we increase our sample size (the number of random permutations).
+2.3. MinHashLSH for Text Deduplication
 The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document.
-However, the need to compare every pair of documents for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
+However, the need to compare every pair of MinHash signatures for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
 Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs.
-MinHashLSH (leskovec2014finding) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents.
+MinHashLSH (Leskovec et al., 2014) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents.
 Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document.
@@
 With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$.
-This procedure enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity.
-This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
-One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication.
+This procedure enables us to produce a series of candidate pairs of documents that we can either take as duplicate pairs (our approach in this work) or further evaluate for similarity.
+This method avoids pairwise comparison between every pair of documents in the corpus but introduces errors due to the use of locality-sensitive hashing.
+One can implement a prefix-tree or a hashmap to search efficiently for candidate duplicate documents when evaluating a new document for duplication.
 However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 277 TB for 5B documents.
-Additionally, inserting and querying these tree-structured indices constitutes a significant portion (¿85%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1).
-Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
-Figure 1. Breakdown of wall clock time on 10% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
+Moreover, navigating these data structures may require pointer-chasing and, therefore, non-contiguous memory accesses that cause cache misses.
+Inserting into and querying against these indices constitutes a significant portion (¿85%) of algorithm runtime, acting as the primary bottleneck for real-time stream throughput at extreme scales (see Figure 1).
+Here, we remedy this space and throughput issue by employing a novel combination of MinHashLSH and Bloom filters.
+Figure 1. Breakdown of wall clock time on 10% of peS2o for conventional MinHashLSH and our LSHBloom method
 2.4. Bloom Filter
-The Bloom filter (bloom1970space) is a space-efficient probabilistic data structure used to test whether an element is a member of a set.
+The Bloom filter (Bloom, 1970) is a space-efficient probabilistic data structure used to test whether an element is a member of a set.
 It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error.
 That is, the Bloom filter may admit false positives but not false negatives.
-A Bloom filter consists of a bit array of size $m$, with all bits initially set to 0.
+A Bloom filter consists of a contiguous bit array of size $m$, with all bits initially set to 0.
 $k$ independent hash functions are defined, each of which hashes an element to an integer in the range 1 to $m$.
@@
 If any of those $k$ bits is 0, the element is not contained in the set.
-Bloom filters have been used extensively in distributed computing for a variety of purposes (broder2004network; chervenak2008globus; tarkoma2011theory).
+Bloom filters have been used extensively in distributed computing for a variety of purposes (Broder and Mitzenmacher, 2004; Chervenak et al., 2008; Tarkoma et al., 2011).
 3. Related Work
-Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs.
-LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data.
-Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 (touvron_llama_2023-1), Claude, Gemini (gemini_team_gemini_2024; gemini_team_gemini_2024-1), and GPT versions 3 and later (openai_gpt-4_2024).
-Others, such as Palm2 (anil2023palm) and Gemma (gemma_team_gemma_2024; gemini_team_gemini_2024-1), provide only limited details.
-Others, including several leading models, provide extensive details about how data for their model was deduplicated including Olmo’s Dolma dataset (soldaini_dolma_2024) as well as its successor DataComp-LM (li2024datacomp), Llama 1 (touvron_llama_2023), Llama 3 (dubey_llama_2024), and OPT (zhang_opt_2022).
-For example, Llama 3 (dubey_llama_2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data.
-Phi uses an approach described by gunasekar_textbooks_2023 and elaborated on in their Phi 2.0 report (javaheripi_surprising_2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation.
-C4 (t5_c4_raffel2020) applies only heuristic filtering for data quality (e.g., removing lines without punctuation or fewer than three words).
-RedPajama (together2023redpajama) uses both exact matching via Bloom filters and fuzzy deduplication via MinHashLSH.
-Datasets such as the Pile (gao2020pile800gbdatasetdiverse) and more recent releases such as FineWeb2 (penedo2025fineweb2pipelinescale) apply MinHashLSH for fuzzy deduplication.
-The most common methods used across these model sources include metadata-based approaches, MinHashLSH, CCNet, and n-gram-based methods.
-Metadata-based approaches (e.g., exact URL matching, DOIs) are by far the simplest approach; however, they are applicable only in cases in which a reliable source of metadata exists, such as those provided by a single publisher.
-Unfortunately, for our target datasets this metadata is often unavailable or is otherwise untrustworthy.
-This can occur, for example, when: 1) a document appears in two or more collections with different metadata schemes (e.g., ArXiV vs.
+The task of streaming text deduplication lies at the intersection of the literature on Set-Similarity joins, Approximate Membership Queries, and Streaming Entity Resolution.
+While traditional approaches prioritize fidelity at the cost of computational resources, LSHBloom targets the Pareto-optimal frontier of high-dimensional streaming deduplication—offering state-of-the-art deduplication fidelity while minimizing resource usage.
+3.1. Similarity Joins
+Traditional techniques for performing a Set-Similarity Join attempt to identify all pairs of records such that their set-similarity (as measured by the Jaccard, Dice, or Cosine metrics) exceeds a given threshold.
+Mann et al.
+(2016) perform a comprehensive evaluation of techniques for performing the Set-Similarity Join (e.g., AllPairs, PPJoin).
+While employing certain filter optimizations can improve upon the quadratic cost of computing exact set-similarity metrics, ultimately, these techniques are not appropriate for extremely high-dimensional entities such as documents.
+Techniques such as SemDedup (Abbas et al., 2023) aim to deduplicate text entries based on their overlap in terms of semantic content rather than using purely textual overlap to compute similarity; however, due to the use of language models, this technique incurs significant computational costs.
+Metadata-based approaches (e.g., exact URL matching, matching DOIs), including blocking using common field values (Christen, 2012), are a simple and efficient alternative, but are applicable only when a reliable source of metadata exists.
+Unfortunately, for our target datasets metadata is often unavailable or is untrustworthy, as when, for example: 1) a document appears in two or more collections with different metadata schemes (e.g., ArXiV vs.
 ACM, different GitHub repositories); 2) sections/paragraphs of text are copied, as in extended journal versions of computer science papers; or 3) metadata is not curated (e.g., an ArXiV entry may not be updated to reflect a subsequent publication).
-CCNet (wenzek2020ccnet) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content.
+3.2. Approximate Membership Queries
+Rather than attempting to directly identify pairs of similar entities, some techniques instead attempt to simply determine, online, whether a new record already belongs to the existing dataset at the time of ingestion.
+For example, by using a special data structure such as a hashmap or Bloom filter (Bloom, 1970) to query for matches against previously ingested records.
+While naive exact-matching with hashmaps or Bloom filters is efficient, it is not robust to the minor variations in text content that are inherent when using text data obtained from the web or parsed from various formats.
+For text, it is preferable to use approximate notions of similarity.
+For instance, the Streaming Quotient Filter (SQF) (Dutta et al., 2013) is an approximate duplicate detection approach for streams of incoming data.
+However, the SQF is optimized for approximate duplicate detection in low-dimensional spaces and is not appropriate for text deduplication.
+For text, techniques such as MinHashing (Broder et al., 1998) and MinHashLSH (Leskovec et al., 2014) are more appropriate due to their ability to compress high-dimensional text records into fixed-length representations.
+The use of LSH functions in MinHashLSH has the additional advantage of identifying candidate duplicates using hashing as opposed to running comparisons with each existing record in the dataset.
+MinHash-based techniques use the Jaccard similarity to measure text overlap; however, n-gram overlap is also a reasonable metric—albeit one that is sensitive to the frequency of n-grams in a given document.
+Frequently, techniques evaluate the n-gram overlap of a document with the n-gram content of the remainder of the corpus using a Bloom filter, trading some granularity of comparison for space efficiency and increased throughput.
+This approach was utilized in preparing the DOLMA (Soldaini and others, 2024) dataset and the DataComp-LM (Li et al., 2024) baseline dataset.
+LSHBloom falls in this category of techniques for online approximate membership queries.
+3.3. State-of-The-Art Deduplication Techniques for Text
+Below, we describe in detail popular techniques for identifying duplicates in text datasets, commonly applied in real-world text dataset preparation.
+CCNet (Wenzek et al., 2020) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content.
+CCNet was employed as part of the data processing workflow for producing the Llama series of models (Touvron et al., 2023; Touvron and others, 2023; Dubey and others, 2024).
 CCNet performs dataset sharding, deduplication, language identification, and quality filtering.
-CCNet’s deduplication technique begins by preprocessing text so that all characters are lowercase and eliminates special Unicode characters.
+CCNet’s deduplication process begins by preprocessing text so that all characters are lowercase and eliminating special Unicode characters.
 CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates.
-CCNet can optionally compare to a single document, a subset of documents, or all documents.
-Dolma (soldaini_dolma_2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma).
-An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter.
+CCNet is only capable of identifying documents that are exactly identical to one another.
+Dolma (Soldaini and others, 2024) deduplicates paragraphs by exact matching using a Bloom filter.
+The Dolma-Ngram alternative splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter.
 If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate.
 By default, $T$ = 1.0 in their configuration.
-DataComp-LM (li2024datacomp) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter.
-They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer (ye2023uniseg).
-For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams.
+These two techniques were used to produce the Dolma dataset (Soldaini and others, 2024) used to produce the Olmo series of language models (Groeneveld et al., 2024; OLMo et al., 2024).
+DataComp-LM (Li et al., 2024) (DCLM) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter.
+This technique was applied to create the DataComp-LM baseline dataset.
+DCLM splits each document into paragraphs on newline characters and tokenizes paragraphs with the UniSeg tokenizer (Ye et al., 2023).
+For each paragraph, it queries each n-gram against its Bloom filter and removes duplicated n-grams.
 If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document.
@@
 They employ the same threshold parameters for paragraph-level and document-level deduplication.
-A table comparing various methods for data deduplication in contemporary LLM pre-training pipelines can be found in Appendix A.
-4. The LSHBloom Method
-4.1. Overview
+MinHashLSH (Leskovec et al., 2014) the MinHashLSH method as described in Section 2.3.
+MinHash-based deduplication was applied in the data ingestion pipeline for OpenAI’s GPT-3 (Brown et al., 2020), DeepMind’s Gopher (Rae et al., 2021), AllenAI’s Olmo3 (Olmo et al., 2025), and Meta’s Llama3 (Dubey and others, 2024).
+MinHash-based deduplication and MinHashLSH were also employed in creating many popular language modeling datasets, such as RedPajama (Computer, 2023), the Pile (Gao et al., 2020) (used for GPT-J (Mireshghallah et al., 2023), GPT-NeoX (Black et al., 2022), and Pythia (Biderman et al., 2023)), FineWeb2 (Penedo et al., 2025), RefinedWeb (Penedo et al., 2023) (used for the Falcon series of models), StarCoder (Li et al., 2023), and StarCoder2 (Lozhkov et al., 2024).
+The sheer number of LLM data pipelines that employ MinHashLSH is a testament to its effectiveness.
+Furthermore, it highlights the potential impact of enhancing its scalability for internet-scale datasets through innovative techniques, such as LSHBloom.
+4. Our Method: LSHBloom
 We propose a memory-efficient alternative to MinHashLSH for document deduplication.
-In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH.
+Rather than employ the traditional prefix-tree or hashmap index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH.
 We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix.
-By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates.
-The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of bender2022optimal.
+By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in duplicate searches.
+The size of each Bloom filter and the number of hash functions utilized are chosen according to the method determined by Bender et al.
+(2022), which guarantees optimal (minimal) space usage.
+LSHBloom effectively acts as a Similarity-Aware Approximate Membership Query data structure, combining the compressive properties of MinHashLSH with the probabilistic indexing of Bloom filters to create a high throughput, space-efficient, online deduplication method for text ingestion workflows.
 4.1. Insertion
@@
 To avoid storing the signatures in full, as is required by MinHashLSH, we assign each band of the minhash signature matrix its own Bloom filter.
-Thus we have in total $b$ Bloom filters, and a collision in any one indicates a duplicate document.
-When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function).
+Thus, we have in total $b$ Bloom filters, and a collision in any one indicates a duplicate document.
+To insert a document into the index, we first compute its MinHash signature, which is a vector of integers (each produced by a universal hash function).
 Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm.
 It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions.
-To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (sha1)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values.
-Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (carter1977universal):
+To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (3rd and Jones, 2001)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values.
+Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (Carter and Wegman, 1977):
 $h(\bar{x})=\left(\sum_{i=1}^{r}h_{i}(x_{i})\right)\mathrm{mod}\ {N}$
@@
 Querying a document against the index is a similar process to insertion.
-We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in subsection 4.2.
+We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in Section 4.1.
 If there is a collision for any band of signatures, we mark that document as a duplicate.
@@
 However, the false positive probability is configurable and can be set to be arbitrarily low.
+An additional benefit of LSHBloom is that the false positive overhead is analytically boundable, as we demonstrate below.
 4.3. Error Rate
 In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate.
-This is also the case in LSHBloom.
-However, we make use of Bloom filters, which incur additional false positive overhead.
+This is also the case in LSHBloom, where the use of Bloom filters incurs additional false positive overhead.
 Here, we analyze the error rate of LSHBloom.
-A false positive is a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom, and a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
+We define a false positive as a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom; a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
 Given a Minhash signature matrix with $b$ bands with $r$ rows each corresponding to $b$ Bloom filters with their individual false positive rates set to $p$, the net false positive overhead of using Bloom filters will be $p\textsubscript{effective}=1-(1-p)^{b}$.
@@
 Therefore, the probability that no Bloom filter reports a false positive is $(1-p)^{b}$, and the overall false positive rate overhead introduced by using Bloom filters is the complement of this value.
+This means we can enforce a false positive overhead of $p\textsubscript{effective}$ by setting the value of $p$ for each Bloom filter as $p=1-(1-p\textsubscript{effective})^{1/b}$
 The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows.
-leskovec2014finding determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
+Leskovec et al.
+(2014) determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
 (1) $FP_{lsh}=\int_{0}^{T}1-(1-t^{r})^{b}dt$
 (2) $FN_{lsh}=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$
-zhu_lshensemble demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
+Zhu et al. (2016) describe how to tune $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
 A Bloom filter will not yield false negatives, but may yield false positives.
 If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result.
-If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$.
-Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.2) for insertion into the Bloom filter.
+If MinHashLSH does not yield a false positive, the Bloom filter index may still yield a false positive result with probability ppeffective.
+Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.1) for insertion into the Bloom filter.
 Therefore, the overall false positive rate for LSHBloom is given by:
-(3) $FP_{bloom}=FP_{lsh}+(1-FP_{lsh})*(p_{effective}+b/N)$
-Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible.
-For 32-bit hashvalues (as is the default in packages such as datasketch) $N=2^{32}$, so the probability of collision for bands prior to insertion into the Bloom filter is also quite small.
-Additionally, because the size of the Bloom filter index scales with the number of items inserted and is not influenced by the size of the integers inserted into it, one can increase the size of the hashvalues (to say, a 64 or 128-bit integer) without increasing the size of the Bloom filter index.
-This is not true of the traditional MinHashLSH index which must store these hashvalues for comparison.
-Meanwhile, for each false negative that MinHashLSH produces, $p_{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the bloom filter index and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion.
+(3) $FP_{\textrm{bloom}}=FP_{lsh}+(1-FP_{lsh})*(p_{\textrm{effective}}+b/N)$
+Note that $p\textsubscript{effective}$ can be set arbitrarily small in practice so that this overhead is negligible.
+For 32-bit hashvalues (the default in packages such as datasketch) $N=2^{32}$, so the collision probability for bands before insertion into the Bloom filter is also small.
+Meanwhile, for each false negative that MinHashLSH produces, $p\textsubscript{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the Bloom filter index, and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion.
 That is,
-(4) $FN_{bloom}=(1-(p_{effective}+b/N))*FN_{lsh}$
+(4) $FN_{\textrm{bloom}}=(1-(p_{\textrm{effective}}+b/N))*FN_{lsh}$
+4.4. Optimizations
+We codesign our implementation of LSHBloom with supercomputer hardware to maximize gains in throughput. We describe these optimizations in detail below.
+4.4.1. Hashing
+After profiling the original implementation of LSHBloom, we optimized the hashing operations used in LSHBloom with careful hardware and software co-design.
+Our profiling revealed that hashing of the integer vectors described in Section 4.1 accounted for over 90% of the time spent in the insert and query routines for LSHBloom.
+The key reason for this was the use of an inefficient software extended-precision integer representation included in Python, which stores extended integers as base-10 strings (Python Software Foundation, 2024).
+By contrast, NumPy offers fixed-precision, C-style integer representations as well as support for vectorized additions, which enhance throughput (Harris et al., 2020; NumPy Developers, 2024).
+However, adding hundreds of fixed-precision integers may lead to overflows, inflating the collision probability of hashvalues before insertion into our Bloom filters.
+The hash of each MinHash permutation is 64 bits, so adding them requires at most 71 bits of unsigned precision to perform without overflow.
+With this in mind, 128-bit arithmetic can be used to avoid integer overflow.
+Modern Intel and AMD x86_64 CPUs implement 128-bit integer arithmetic using the adc (Guide, 2011) (”add with carry“) instruction with equivalents such as addc.u64 on Nvidia PTX (Nickolls et al., 2008; NVIDIA Corporation, 2024) on other platforms.
+Additionally, the modular division operation performed as the last step can be implemented either using a simple bit mask or by simply using the lower order half of the 128-bit register resulting from the add with carry operation.
+Using Rust’s standard extended precision support (Matsakis and Klock, 2014; The Rust Project Developers, 2024) makes expressing this operation easy and results in only 30 lines of optimized x86_64 assembly depending on architecture and loop unrolling decisions.
+We replace our Python logic for the integer vector hashing operation described in Section 4.1 with this Rust function.
+This single optimized function is over $94\%$ faster than the existing Python implementation, resulting in a $11\times$ end-to-end wall clock improvement to LSHBloom.
+4.4.2. Parallelization and Shared Memory
+Parallelization is an obvious strategy for acceleration.
+As shown in Figure 1, MinHashing accounts for the majority of the runtime in LSHBloom.
+Thankfully, each document can be MinHashed independently and this step can easily be parallelized across multiple nodes using frameworks such as Parsl (Babuji et al., 2019).
+On single-node runs (as in our evaluation), we can employ multiprocessing to speed up this step.
+When finally writing to the Bloom filter index, we do so sequentially (without parallelism).
+This is because, at the time of insertion, we must also ensure that the current document is not a duplicate of a previously inserted one.
+Naively parallelizing this step would lead to errors in duplicate identification.
+Parallelizing regardless and insisting upon synchronizing Bloom filters across threads or nodes would introduce significant communication overhead.
+Despite this lack of parallelism, our optimized hashing routines still enable extremely fast insert and query operations.
+Furthermore, we take advantage of the tiny size of our Bloom filter indices by hosting our Bloom filters in node-local shared memory segments (via /dev/shm), allowing us to locate our index in DRAM with swap partitions on local SSDs.
+This further accelerates insert and query operations on the Bloom filter index.
 4.5. Advantages
-We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams.
-Since many practical scenarios may admit an extremely large universe of such n-grams (e.g., if we allow our set of documents to include a variety of encodings, languages, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the size of MinHash hashvalues (to 64 or 128-bits or more) or increasing the accuracy of MinHash estimates of the Jaccard similarity by increasing the number of MinHash permutations.
-A scalable deduplication algorithm should be amenable to such considerations.
-LSHBloom allows for the space-efficient representation of extremely large deduplication indices.
-This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with a fixed number of bits.
-This means that increasing the size of hashvalues does not change the index size.
-Furthermore, using more MinHash permutations only increases the index size linearly with the number of bands that are required for the MinHash signature matrix–which is far fewer than the number of permutations.
-If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (bender2022optimal).
+LSHBloom allows for the space-efficient deduplication of extremely large corpora.
+This is due to the Bloom filter’s ability to determine the set membership properties of objects (e.g., documents) of arbitrary size with a fixed number of bits.
+The size of the LSHBloom index only depends on the number of planned documents, $n$, the number of bands, $b$, in the MinHash signature matrix, and the desired false positive rate bound, $p$, for each Bloom filter.
+If we expect $n$ documents and desire a false positive bound of $p$, then each of our $b$ Bloom filters requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (Bender et al., 2022).
+This means that the size of our index is invariant to the size of MinHash hashvalues—we can increase the size of MinHash hashvalues (to say, a 64 or 128-bit integer) to enhance the accuracy of our Jaccard similarity estimate at no additional cost.
+Meanwhile, because the traditional MinHashLSH index must store these hashvalues for comparison, increasing the size of the MinHash hashvalues results in a linear increase in the size of the traditional LSHIndex.
+Enhancing the accuracy of our probabilistic estimate of the Jaccard similarity by increasing the number of MinHash permutations we apply will increase the size of the MinHash signature matrix, which will only change the size of the Bloom filter index if the number of bands, $b$, in the signature matrix increases.
+Luckily, the number of bands in the signature matrix is generally far fewer than the number of total permutations.
 As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing.
-For these parameter settings, MinHashLSH creates nine bands (using the method described in (zhu_lshensemble) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index.
+For these parameter settings, MinHashLSH creates nine bands (using the method described in (Zhu et al., 2016) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index.
 For these settings, setting a conservative ppeffective = 1e-10 for a dataset of 10B documents would only require 590 GB of space.
-Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex—exemplifying the scalability of our technique.
-While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration.
-This is another advantage of our technique: we can observe massive gains in space while only introducing a single additional hyperparameter, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
-The additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter.
-Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter.
-However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.
+Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex.
+While the use of Bloom filters to satisfy memory and throughput constraints introduces additional false positive errors, the rate of false positives and false negatives is analytically boundable.
+Moreover, the additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter.
+Therefore, there is a trade-off between space efficiency and correctness that is mediated by the size of each filter.
+However, as demonstrated by our example above, this trade-off tends to be highly favorable even on extremely large datasets.
+Additionally, LSHBloom only introduces a single additional hyperparameter, the false positive rate for our Bloom filters, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
+Beyond storage savings, the Bloom filter index yields substantial enhancements in ingestion throughput due to two primary factors.
+1) Computational efficiency: insertion and querying against the index requires only a few independent hash operations and modular arithmetic, which is fast on modern hardware (see Section 4.4), and 2) Cache-Aware Data layout: in contrast to the pointer chasing inherent in tree or chained-hashmap-based indices, LSHBloom’s use of contiguously allocated bitarrays maximizes spatial locality and minimizes cache misses.
 5. Evaluation
-Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance.
-As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets.
-To assess correctness, we tune LSHBloom and other state-of-the-art deduplication methods on a balanced dataset of 25K PDFs containing labeled duplicates.
-We then use the best parameter settings for each algorithm and compare them on a larger labelled dataset of 50K PDFs.
-Without sufficient correctness, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary training overhead; or over-identifying duplicates, which unnecessarily reduces the training dataset volume.
-For performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (peS2o).
-Without sufficient performance, it may not be possible to process large datasets on available hardware resources.
-In the following subsections, we describe the details of our evaluation beginning with the evaluation methodology, then proceeding to correctness, performance, and finally the tradeoffs between performance and correctness.
+In evaluating a deduplication method for a particular application, we are concerned with its deduplication fidelity, throughput, and disk space requirements.
+We measure deduplication fidelity via the precision and recall metrics.
+Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces dataset size and, in deep learning, limits an LLM’s ability to learn the underlying distribution of the text.
+Low recall, i.e., excessive false negatives, admits a great degree of duplicate content, with negative effects on LLM quality in pre-training applications.
+High throughput and a small disk footprint is critical due to the extreme size of modern LLM datasets.
+For example, peS2o (Soldaini and Lo, 2023) contains roughly 39 million documents, DOLMA (Soldaini and others, 2024) 5B documents, SmolLM3’s pretraining dataset has over 15B documents (Team, 2025), and RedPajama (Computer, 2023) has over 100B documents—scales at which naive pairwise comparisons are infeasible.
+Therefore, deduplicating at the billion-record scale with high fidelity is paramount.
+To assess deduplication fidelity, we tune LSHBloom and other state-of-the-art deduplication methods on a synthetic, balanced dataset of 24K PDFs containing labeled duplicates, we refer to this as our tuning dataset.
+We then use the best parameter settings for each algorithm and compare them on a larger synthetic dataset of 50K PDFs which we term the testing dataset.
+Without sufficient fidelity, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary processing overhead, or over-identifying duplicates, which unnecessarily reduces training dataset volumes.
+To measure performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (Soldaini and Lo, 2023).
+Without a sufficiently high throughput and sufficiently low disk footprint, it may not be possible to process large datasets on available hardware resources.
+In the following subsections, we describe the details of our evaluation, beginning with the evaluation methodology, then proceeding to our evaluation of fidelity, performance, and finally the tradeoffs between performance and fidelity.
 5.1. Experimental Methodology
-In this subsection, we describe our experimental methodology including testbed hardware, state-of-the-art methods for comparison, our evaluation metrics, and datasets.
+We describe our experimental methodology, including testbed hardware, state-of-the-art methods against which we compare, evaluation metrics, and datasets.
 5.1.1. Testbed Hardware
@@
 5.1.2. Baselines
-We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM.
-We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language.
-This also enabled us to normalize the Bloom filter implementation for those techniques which use them.
+We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH (Broder, 1997), Dolma and Dolma-Ngram (Soldaini and others, 2024), CCNet (Wenzek et al., 2020), and DataComp-LM (Li et al., 2024) (see 3.3 for details).
+We normalized all implementations to Python implementations to avoid performance differences for common procedures due to implementation language.
+This also enabled us to normalize the Bloom filter modules for those techniques that use them to a fast C/Python-based implementation provided by pybloomfiltermmap3 (Sinha, 2025).
+LSHBloom is evaluated using the optimized hashing routine described in Section 4.4.
 Below, we describe any changes or configurations to the methods.
-MinhashLSH (broder1997resemblance) (lsh): We configure the traditional MinhashLSH algorithm as described in subsection 2.3.
+MinhashLSH (Broder, 1997): We configure traditional MinhashLSH as described in Section 2.3.
+We use the implementation offered by the popular datasketch package (Zhu, 2025), which implements a hashmap-based LSHIndex.
+We compute MinHashes using multiprocessing for both MinHashLSH and LSHBloom.
 DOLMA: We apply both standard Dolma and Dolma-Ngram methods to document-level deduplication.
-Specifically, for dolma we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$.
-For dolma-ngram, we split each document into a set of n-grams with a given size, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams that are duplicated within that document to the overlap threshold, $T$.
+For Dolma, we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$.
+For Dolma-Ngram, we split each document into n-grams, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams duplicated within that document to the overlap threshold, $T$.
 Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication.
-CCNet originally performs deduplication at the paragraph level; however, here, we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1].
-For a fair comparison, we compare documents to all other documents—just as MinhashLSH does.
-When evaluating memory usage, we apply all CCNet optimizations, such as the use of memory-efficient hash sets.
-Additionally, we augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a more appropriate comparison during our scaling tests.
+CCNet originally performs deduplication at the paragraph level; we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1].
+We augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a fair comparison during our scaling tests.
 DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
-When considering the correctness of techniques that operate on n-grams and use Bloom filters, we must give an expected n-gram count in the corpus; choosing this count accurately is crucial to make a fair comparison.
-Because counting the number of n-grams of a given size in a corpus is far more computationally expensive than counting the number of documents (as is required for our technique), we compute an estimate of the n-gram count in a corpus.
-To do so we randomly sample $N$ = 1000 documents from the corpus, compute the mean n-gram count from our sample, and multiply the mean n-gram count by the number of total documents in that corpus to obtain an estimate of the number of n-grams in the entire corpus.
-We parameterize Bloom filters for n-gram techniques using this estimate.
-We apply an analogous procedure to obtain estimated paragraph counts for Dolma and CCNet.
+For Dolma, Dolma-Ngram, and DCLM, we must give an expected n-gram/paragraph count for the corpus; choosing this count accurately is crucial to making a fair comparison because it influences Bloom filter size and error rates.
+Because counting the number of n-grams/paragraphs in a corpus is very expensive compared to counting documents (as is required for MinHash-based techniques), we compute an estimate of the n-gram/paragraph count in a corpus by randomly sampling $N$ = 1000 documents from the corpus, computing the mean n-gram/paragraph count from our sample, and multiplying the mean count by the number of total documents in that corpus to obtain an estimated total n-gram/paragraph count.
 5.1.3. Metrics
-Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents.
+Fidelity: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents.
 Here, we consider a ”positive” label to mean that a document is a duplicate of another document in the corpus.
-We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance.
+We want these methods to yield few false positives (to retain the maximal number of informative records), few false negatives (to reduce the prevalence of duplication), and to provide a reasonable combination of accuracy and performance.
 Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives.
-Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization.
+Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and decreases model quality.
 To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using its precision (the proportion of duplicate labels that are true duplicates), its recall (the proportion of duplicates in the dataset that we correctly identify), and its F1 score (the harmonic mean of precision and recall):
 $F1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$
-Performance: To assess the resource costs of each technique we measure wall clock time and disk space usage.
-We measure wall clock time using /user/bin/time and assess each method’s disk space requirements using Python’s standard os.path.getsize method.
-5.1.4. Constructing the correctness evaluation dataset
+Performance: To assess the resource costs of each technique, we measure wall clock time (using /user/bin/time) and disk space requirements (using Python’s standard os.path.getsize).
+5.1.4. Constructing the fidelity evaluation dataset
 To permit the rigorous evaluation of de-duplication methods, we created a pipeline for generating labeled datasets of duplicate documents.
-We acquire scientific articles (scraped from HTML and parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata.
+Siebenschuh et al.
+(2025) produced a dataset of parsed PDFs for evaluating their adaptive PDF parsing technique, AdaParse.
+We utilize this baseline dataset to construct synthetic tuning and testing datasets for the deduplication task.
+The baseline dataset consists of scientific articles with two versions: one scraped from HTML and the other parsed from PDF.
+These span diverse disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata.
 Text content from HTML is directly extracted.
-We translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF (PyMuPDF), Nougat (blecher2023nougat), or Tesseract (smith2007overview), each used with roughly the same frequency.
-We then can construct datasets where documents are duplicated, either by including multiple versions of the same document which were parsed using different text extraction tools and/or by including randomly truncated versions of that document.
-This reflects common forms of duplication that occur in real-world web scraping pipelines where the same document can be ingested via different text extraction tooling either due to differences in the origin of the document instance, the fact that the same document is often found in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors which can abruptly skip or truncate portions of the document’s text.
-We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm.
-We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances.
+Each PDF is parsed by applying a text extraction or OCR tool—one of PyMupDF (Smith, 2024), Nougat (Blecher et al., 2023), or Tesseract (Smith, 2007), each used with roughly the same frequency.
+We construct datasets containing duplicated documents from this collection of parsed texts, either by including multiple versions of the same document parsed with different tools and/or by including randomly truncated versions of that document.
+We balance the number of parser and truncation-based duplicates in each dataset we construct, which prevents evaluation bias towards techniques that are better suited to identifying just one of these types of duplication.
+These scenarios reflect common forms of duplication encountered in real-world web scraping pipelines where the same document may be ingested via different text extraction tooling due to differences in document origin, the same document being obtained in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors, which can abruptly skip or truncate portions of the document’s text.
+We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm, the tuning dataset.
+We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances, the testing datasets.
 5.1.5. Choice of values for Tuning Parameters
-We evaluate each technique over a range of evenly spaced parameter settings using a grid search.
-Specifically, for any threshold parameters such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values between 0.2 and 1.0 in steps of 0.2.
-We vary the number of permutations for MinHashLSH from 32 to 256 by powers of two.
+We evaluate each technique for a range of parameter settings.
+Specifically, for each threshold parameter, such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values from 0.2 to 1.0 in steps of 0.2.
+We vary the number of permutations for MinHashing from 32 to 256 by powers of two.
 For n-gram techniques, we explore n-gram sizes 1, 2, 5, 7, 13, and 26.
-For MinHashLSH and LSHBloom, we additionally gathered results for $T=0.5$ and 48 random permutations as the surrounding parameter values demonstrated sharp jumps in the F1 score, and we wanted to explore this change in the score at a finer granularity.
-All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM, LSHBloom) are assigned a false positive rate $p$ of 1e-5.
-For LSHBloom, this equates to $p\textsubscript{effective}$ between 1e-5 and 8e-4, depending on the value of $T$ and the number of MinHash permutations.
-The value of $p_{effective}$ at various parameter settings for this benchmark is illustrated in Figure 2.
-Figure 2. Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given NN=24,956 documents and pp=1e-5
+For MinHashLSH and LSHBloom, we also gathered results for $T$ = 0.5 and 48 random permutations, as the surrounding parameter values demonstrated sharp jumps in F1 score, and we wanted to explore this change in the score at a finer granularity.
+All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM) are assigned a false positive rate $p$ of 1e-5.
+For LSHBloom, we set $p\textsubscript{effective}$ to 1e-5.
 It is possible that one could properly tune these deduplication methods further using more incremental changes in their parameter settings.
-However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance.
+However, the expensive runtime of n-gram techniques on large corpora, coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance.
 Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs.
-Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations.
+Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations (Leskovec et al., 2014).
 MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget.
-This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see subsubsection 5.4.1).
-(a) MinHashLSH
-(b) LSHBloom
-Figure 3. F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
-5.2. Hyperparameter Tuning Results (24K Dataset)
+This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see Section 5.4.1).
+Figure 2. F1 score for LSH techniques as a function of number of permutations (x-axis) and Jaccard similarity threshold (y axis).
+5.2. Hyperparameter Tuning
 5.2.1. LSH Methods
-The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts.
-Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 24K dataset.
-As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (broder1998min).
+The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects on their accuracy, so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts.
+Figure 2 shows the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the tuning dataset.
+As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (Broder et al., 1998).
 Correspondingly, we use 256 MinHash permutations in our subsequent experiments.
@@
 For our particular benchmark, a threshold of 0.5 yields the best performance; more stringent thresholds lead to a diminishing F1 score.
-(a) DOLMA
-(b) DataComp-LM
 Figure 3. F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
 5.2.2. Paragraph Level and N-Gram Methods
-Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score.
-Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance.
+Figure 3 shows the results of varying both the n-gram size and overlap threshold parameters for each n-gram method.
+DCLM is fairly successful, even approaching the F1 score of the MinHashLSH-based techniques, and demonstrates moderate sensitivity to the overlap threshold parameter.
+Smaller n-gram sizes seem to perform better as well—likely due to the increased granularity with which they enable comparison between a document and the remainder of the corpus.
+For DCLM, the best setting is using an overlap threshold of 0.2 and an n-gram size of 5.
+In contrast, Dolma-Ngram seems to be fairly insensitive to the choice of parameters and performs relatively poorly.
+The difference between these two very similar techniques may stem from DCLM’s use of the UniSeg tokenizer, which splits text into meaningful units based on rules defined by the Unicode standard—Dolma-Ngram simply splits text by whitespace.
+Figure 4 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on F1 score.
+Predictably, distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance.
 The best F1 for paragraph-level methods is achieved by setting a low threshold of 0.2.
-N-gram methods such as DCLM do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques.
-Interestingly, DOLMA and DOLMA-ngram perform very similarly.
-Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method.
-While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus.
-This is similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error.
+While n-gram methods enable finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus.
+This is somewhat similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error.
 Moreover, paragraph-level techniques demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions.
@@
 Table 1 shows the parameter settings for each technique that we use for the remainder of our experiments.
-Figure 5. F1 score vs. threshold for paragraph-level techniques.
+Figure 4. F1 score vs. threshold, paragraph-level techniques.
 Table 1. Best settings for each deduplication technique
@@
 CCNet - 0.2
-5.3. Deduplication Correctness (50K Dataset)
-(a) Precision
-(b) Recall
-(c) F1
-Figure 6. Correctness metrics for deduplication techniques for various levels of duplication.
+5.3. Deduplication Fidelity
+Figure 5. Fidelity metrics for deduplication techniques for various levels of duplication.
 5.3.1. Deduplication Quality
-Here we evaluate the correctness of each technique using metrics such as their precision, recall, and F1 score.
+Here we evaluate the fidelity of each technique using metrics such as precision, recall, and F1 score.
 We do so by constructing datasets of 50,000 documents each with varying levels of duplication between 10% and 90%.
-We construct these datasets using the procedure described in subsubsection 5.1.4.
+We construct these datasets using the procedure described in Section 5.1.4.
 Figure 5 illustrates the precision, recall, and F1 score of each technique as we vary the proportion of duplicate documents in our dataset.
-Notably, in terms of F1 score (see 6(c)), both MinHashLSH and LSHBloom outperform all other methods—only being slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority ($60\%$) of the dataset consists of duplicates, which is highly unlikely in real-world scenarios.
-MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match the performance of MinHashLSH while affording considerable gains in scalability (to be shown below in subsection 5.4).
-Looking more closely, 6(a) illustrates that MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset.
-The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision often tracking closely with that of the MinHash-based techniques.
+Notably, in terms of F1 score, both MinHashLSH and LSHBloom outperform all other methods—only slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority (>>60%) of the dataset consists of duplicates, highly unlikely in real-world scenarios.
+MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match MinHashLSH’s performance while affording considerable gains in scalability (to be shown below in Section 5.4).
+Looking more closely, MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset.
+The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision tracking closely with that of the MinHash-based techniques.
 However, Dolma’s dismal recall scores (<0.4<0.4) suggest that this is likely the result of under-identifying duplicate instances rather than being a reflection of greater capacity for discernment.
@@
 This is due to a greater number of true positives being present in the dataset.
-In terms of recall, 6(b) demonstrates that DCLM and Dolma-Ngram tend to perform the best—though notably their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare.
+In terms of recall, DCLM and Dolma-Ngram tend to perform the best, although their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare.
 This indicates that, while these techniques identify most duplicates, they struggle to differentiate between duplicate and non-duplicate entities.
@@
 Our experimental results indicate that LSHBloom achieves deduplication performance comparable to the state-of-the-art—indeed, LSHBloom achieves essentially identical performance to MinHashLSH, one of the most popular choices for text deduplication, with only a marginal degradation in precision.
+(a) Quality vs. Wall Clock Time
+(b) Quality vs. Disk
+Figure 6. Pareto plots for F1 score vs. resource usage.
 5.3.2. Performance vs. Quality Tradeoffs
@@
 We visualize data from our evaluation of the balanced dataset of 50,000 documents.
-Figures 7(a) and 7(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively.
+Figures 6(a) and 6(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively.
 MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime.
-LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see subsection 5.4).
+LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see Section 5.4).
 LSHBloom also displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index.
-As shown below in subsection 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
+As shown below in Section 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
 Our results indicate that LSHBloom intelligently discerns between duplicate and non-duplicate documents (as demonstrated by its high precision scores) and can adeptly identify the majority of duplicate documents (as demonstrated by its strong recall scores).
@@
 Overall, however, both yield similar F1 scores.
-This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication use-cases.
+This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication.
 5.4. Deduplication Resource Usage at Scale
-(a) Quality vs. Wall Clock Time
-(b) Quality vs. Disk
-Figure 7. Pareto plots for F1 score vs. resource usage.
 As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents.
-Here we work with the peS2o dataset (peS2o) of roughly 39M documents.
+Here we work with the peS2o dataset (Soldaini and Lo, 2023) of roughly 39M documents.
 We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter).
 We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost.
-DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in subsection 5.3.
+DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in Section 5.3.
 Even processing just 10% of peS2o using this technique would require over 200 hours by extrapolation.
+We find that Dolma-Ngram is similarly prohibitive in runtime, likely due to the time cost of processing many n-grams in documents that are frequently much larger than those in our previous benchmark.
 We discuss the results for other methods below.
-Figure 8. Disk size of index for different methods, as measured for peS2o subsets of increasing size.
-Figure 9. Wall clock times for different methods, as measured for peS2o subsets of increasing size.
+(a) Wall clock times over subsets of peS2o.
+(b) Disk Usage over subsets of peS2o.
+Figure 7. Resource usage at scale for each deduplication technique.
+Figure 8. Extrapolated wall clock times per method at scale.
 5.4.1. Scaling of Resource Usage
@@
 5.4.2. Estimated Resource Usage at Scale
-Figure 10. Predicted wall clock times vs. number of documents (in Billions) for different methods
 […]
 […]
-Table 2. Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
-Technique Bloom filter false Index disk
-positive overhead usage (TB)
-MinHashLSH 277.68
-LSHBloom 1e-5 8.33
-LSHBloom 1e-8 12.11
-LSHBloom 1/5B 15.50
-Table 3. Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
-Technique Bloom filter false Index disk
-positive overhead usage (TB)
-MinHashLSH 555.35
-LSHBloom 1e-5 16.66
-LSHBloom 1e-8 24.21
-LSHBloom 1/100B 31.76
+Table 2. Comparison of Extrapolated Index Storage Requirements
+Technique Bloom FP Index Disk Usage (TB)
+Overhead $N=5\times 10^{9}$ $N=10^{11}$
+MinHashLSH 277.68 555.35
+LSHBloom $10^{-5}$ 8.33 16.66
+$10^{-8}$ 12.11 24.21
+$1/N$ 15.50 31.76
 6. Conclusions
 […]
-In the future, we aim to further reduce the wall clock time of LSHBloom by employing parallelization over subsets of text datasets when inserting them into our index—an approach that we expect to enable significant acceleration.
-Additionally, though very useful, our synthetic benchmark has drawbacks—namely, in its biased and limited notion of what constitutes a proper duplicate.
-Ideally, future work will involve benchmark datasets that cover more diverse forms of duplication in proportion to their representation ”in the wild”—that is, in real-world LLM pre-training data pipelines.
-Appendix A Comparison of Data Deduplication Pipelines for LLM Pre-training
-Table 4. Comparative Analysis of Data Deduplication Pipelines for LLM Pre-training
-Dataset / Model Primary Data Sources Key Cleaning & Filtering Steps Dedup. Level Deduplication Technique(s) Key Parameters & Implementation Notes Citations
-C4 (for T5) Common Crawl (April 2019 snapshot) Heuristic filtering: remove lines without terminal punctuation, lines << 3 words, docs << 5 sentences, docs with ”Lorem ipsum”, and docs containing words from a blocklist.
-Language filtering for English (>>99% probability).
-None Heuristic Filtering Only No explicit deduplication algorithms were used.
-The focus was on rule-based filtering.
-They remove all but one instance of any three-sentence span that occurs more than once in the dataset.
-(t5_c4_raffel2020)
-The Pile (for GPT-Neo) 22 diverse sources including Common Crawl (Pile-CC), Books3, arXiv, GitHub, OpenWebText2, and PubMed Central.
-Source-specific cleaning (e.g., converting TeX to Markdown for arXiv, parsing top answers for StackExchange).
-The entire dataset was filtered for duplicates.
-Document MinHashLSH The OpenWebText2 component specifically used in-memory MinHashLSH via the datasketch library for near-duplicate detection.
-(gao2020pile800gbdatasetdiverse)
-GPT-3 A mix of Common Crawl (filtered), WebText2, Books1, Books2, and Wikipedia.
-Quality filtering was applied to the Common Crawl portion.
-Document MinHashLSH Spark’s MinHashLSH implementation with 10 hashes.
-(brown2020language)
-RedPajama / SlimPajama (LLaMA Replication) LLaMA sources: Common Crawl, C4, GitHub, Books, arXiv, StackExchange.
-Multi-stage pipeline: NFC Unicode normalization, filtering documents << 200 characters.
-Document Exact Hashing (Bloom Filter) & Fuzzy Deduplication (MinHashLSH) RedPajama-V2 uses a Bloom filter for efficient exact deduplication in addition to MinHashLSH.
-SlimPajama uses a distributed MinHashLSH implementation with 13-grams.
-(together2023redpajama; cerebras2023slimpajama)
-CCNet CommonCrawl Language identification (for splitting into monolingual datasets) using lightweight classifiers.
-LM-based quality filtering at the document level.
-Exact (SHA1 Hashing) Line-level Computes SHA1 hashes for each line of text and performs exact deduplication at the line-level using these hash values.
-(wenzek2020ccnet)
-Llama 3 Constructed ”from a variety of data sources containing knowledge until the end of 2023”.
-PII and safety filtering.
-Heuristic filtering: removed lines with repeated content, removed lines with ”dirty” words, filtered out documents with excessive numbers of outlier tokens compared to the rest of the training corpus.
-Document and Line-Level URL (metadata), MinHash, CCNet The paper mentions the use of URL deduplication, MinHash for document-level deduplication, and CCNet for line-level deduplication.
-(dubey_llama_2024)
-Dolma Collected from CommonCrawl, Github, Reddit, Wikipedia, among other sources.
-Remove data containing PII or toxic material.
-Not Specified Paragraph Deduplicates paragraphs via exact-matching (Bloom filter).
-Alternatively, deduplicates paragraphs by measuring their n-gram overlap with the rest of the corpus, also using a Bloom filter.
-Preprocess some web data using CCNet and metadata-based deduplication.
-(soldaini_dolma_2024)
-DataComp-LM (DCLM) Common Crawl (provides a candidate pool of 240T+ tokens).
-A benchmark for data curation strategies.
-The DCLM-Baseline dataset was created using model-based filtering (data quality classifiers) to select high-quality data.
-Paragraph and Document-level Baseline dataset is deduplicated using Bloom filters to measure n-gram overlap of a paragraph/document with the rest of the dataset.
-The DCLM-Baseline dataset of ~4T tokens is combined with StarCoder and ProofPile2.
-(li2024datacomp)
-FineWeb / FineWeb-Edu / FineWeb2 Common Crawl (96 snapshots for FineWeb, ~100 for FineWeb2).
-URL filtering (blocklist), language classification (fastText), and quality/repetition filters.
-FineWeb-Edu uses a classifier trained on Llama-3-70B annotations.
-Document Fuzzy Deduplication (MinHash) Extensive ablation studies.
-The final strategy involved deduplicating each Common Crawl snapshot *individually*, as global deduplication was found to discard high-quality data from older snapshots.
-(penedo2025fineweb2pipelinescale)
-Lee et al.
-(2021) C4, RealNews, LM1B, Wiki-40B-en N/A (focus was on the deduplication technique itself).
-Substring (Exact) Suffix Array Removed all but one occurrence of any 50-gram (sequence of 50 tokens) that appeared in multiple documents.
-This method targets boilerplate text.
-(lee2021deduplicating)
-D4 (SemDeDup) Common Crawl (CC-dedup) An initial pass of MinHash-based deduplication was applied before the main technique.
-Document (Semantic) Semantic Deduplication (Embeddings + K-Means) A multi-stage method that first applies semantic deduplication (SemDeDup) to remove dense clusters of semantically redundant information.
-(tirumala2023d4)
-LSHBloom peS2o (a dataset of 39 million technical documents) N/A (focus was on the deduplication technique itself).
-Document (Approximate) MinHashLSH with Bloom Filter Index An extension of MinHashLSH that replaces the large, memory-intensive LSH index with a series of lightweight Bloom filters (one for each band).
-$12\times$ faster, $18\times$ less space usage.
+In the future, we will further enhance LSHBloom’s throughput by carefully employing parallelization over subsets of text datasets when inserting them into our index.
+Additionally, though very useful, our synthetic benchmark has drawbacks—namely, in that it mainly emulates forms of duplication native to web scraping-based ingestion.
+Ideally, future work will involve benchmark datasets that cover more diverse forms of data duplication.